### v2

# arXiv:2408.13769v2[math.LO] 01 Mar 2026

Abstract
In this article, we try to formulate a definition of many-valued logical structure. For this, we embark on a deeper study of Suszko’s Thesis (ST) and show that the truth or falsity of ST depends, at least, on the precise notion of semantics. We propose two different notions of semantics and three different notions of entailment. The first one helps us formulate a precise definition of inferentially many-valued logical structures. The second and the third help us to generalise Suszko Reduction and provide adequate bivalent semantics for monotonic and a couple of nonmonotonic logical structures. All these lead us to a closer examination of the played by language/metalanguage hierarchy vis-á-vis ST. We conclude that many-valued logical structures can be obtained if the bivalence of all the higher-order metalogics of the logic under consideration is discarded, building formal bridges between the theory of graded consequence and the theory of many-valued logical structures, culminating in generalisations of Suszko’s Thesis.
Keywords: Suszko’s thesis, Universal logic, Semantics, Inferential many-valuedness, Graded consequence, Many-valued logical structures.
1 Introduction
Undoubtedly, a fundamental problem concerning
many-valuedness is to know what it really is.
[14, p. 281]
A logical structure (see [3, p. 84]) is a pair $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$, where $\mathscr{L}$ is a set, $\mathcal{P}(\mathscr{L})$ denotes the power set of $\mathscr{L}$ and $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,\subseteq\,\mathcal{P}(\mathscr{L})\times\mathscr{L}$. Alternatively, a logical structure may be defined as a pair $(\mathscr{L},W)$, where $\mathscr{L}$ is a set and $W:\mathcal{P}(\mathscr{L})\to\mathcal{P}(\mathscr{L})$ is called a consequence operator. These two definitions are equivalent in the following sense.
  1. $\bullet$
    Every logical structure of the form $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$ induces a logical structure of the form $(\mathscr{L},W^{\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}})$ where $W^{\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}}:\mathcal{P}(\mathscr{L})\to\mathcal{P}(\mathscr{L})$ and for all $\alpha\in\mathscr{L}$
    $\alpha\in W^{\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}}(\Gamma)~\text{iff}~\Gamma\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\alpha$
  2. $\bullet$
    Every logical structure of the form $(\mathscr{L},W)$ induces a logical structure of the form $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{W})$ where $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{W}\,\subseteq\,\mathcal{P}(\mathscr{L})\times\mathscr{L}$ and for all $\alpha\in\mathscr{L}$
    $\Gamma\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{W}\alpha~\text{iff}~\alpha\in W(\Gamma)$
Henceforth, we will use either of these descriptions interchangeably.
Although the terms ‘logic’ and ‘logical structure’ have been considered synonymous in [3, p. 84], we distinguish between these here. The term logic is reserved for referring to those pairs $(\mathscr{L},W)$ where $\mathscr{L}$ is an algebra.
Given a set $\mathscr{L}$, a logical matrix for $\mathscr{L}$ is a triple $\langle\mathcal{V},D,S\rangle$ where $D\subseteq\mathcal{V}$, $\mathscr{L}$ and $\mathcal{V}$ are algebras of similar type, and $S$ is the set of all homomorphisms from $\mathscr{L}$ to $\mathcal{V}$. The elements of $S$ are often referred to as valuations. To any such matrix $\mathfrak{M}=\langle\mathcal{V},D,S\rangle$ for $\mathscr{L}$, one can associate a relation $\models_{\mathfrak{M}}\,\subseteq\,\mathcal{P}(\mathscr{L})\times\mathscr{L}$ as follows:
$\Gamma\models_{\mathfrak{M}}\alpha~\text{iff for all}~v\in S,v(\Gamma)\subseteq D~\text{implies that}~v(\alpha)\in D$
for all $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$. This relation is often referred to as the entailment relation induced by the matrix $\mathfrak{M}$. It satisfies the following properties.
  1. (1)
    For all $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$, $\Gamma\models_{\mathfrak{M}}\alpha$. (Reflexivity)
  2. (2)
    For all $\Gamma\cup\Sigma\cup\{\alpha\}\subseteq\mathscr{L}$ such that $\Gamma\subseteq\Sigma$, $\Gamma\models_{\mathfrak{M}}\alpha$ implies that $\Sigma\models_{\mathfrak{M}}\alpha$. (Monotonicity)
  3. (3)
    For all $\Gamma\cup\Sigma\cup\{\alpha\}\subseteq\mathscr{L}$, if $\Gamma\models_{\mathfrak{M}}\beta$ for all $\beta\in\Sigma$, and $\Sigma\models_{\mathfrak{M}}\alpha$ then $\Gamma\models_{\mathfrak{M}}\alpha$ as well. (Transitivity)
In terms of $W^{\models_{\mathfrak{M}}}$, the above properties can be formulated as follows.
  1. (1)
    For all $\Gamma\subseteq\mathscr{L}$, $\Gamma\subseteq W^{\models_{\mathfrak{M}}}(\Gamma)$. (Reflexivity)
  2. (2)
    For all $\Gamma\cup\Sigma\subseteq\mathscr{L}$ such that $\Gamma\subseteq\Sigma$, $W^{\models_{\mathfrak{M}}}(\Gamma)\subseteq W^{\models_{\mathfrak{M}}}(\Sigma)$. (Monotonicity)
  3. (3)
    For all $\Gamma\cup\Sigma\subseteq\mathscr{L}$, $\Gamma\subseteq W^{\models_{\mathfrak{M}}}(\Sigma)$ implies that $W^{\models_{\mathfrak{M}}}(\Gamma)\subseteq W^{\models_{\mathfrak{M}}}(\Sigma)$. (Transitivity)
A logical structure $(\mathscr{L},W)$ for which $W$ satisfies the above three conditions is said to be a Tarski-type logical structure. A logical structure $(\mathscr{L},W)$ is reflexive (respectively, monotonic, transitive) if $W$ satisfies reflexivity (respectively, monotonicity, transitivity).
Suszko’s Thesis ($\mathbf{ST}$), named after the Polish logician Roman Suszko, is the (philosophical) claim that “there are but two logical values, true and false” (see [11, p. 169]). This thesis has been given formal contents by the so-called Suszko Reduction (the particular case of Łukasiewicz’s 3-valued logic $\textbf{Ł}_{3}$ is outlined in [37, pp. 378-379] by Suszko himself). However, as has been pointed out in [10, p. 6], “there seems to be no paper where Suszko explicitly formulates (SR)[the Suszko Reduction] in full generality!” Furthermore, there seems to be no explicit definition of ‘logical value’. This makes $\mathbf{ST}$ rather ambiguous, difficult to justify and open to interpretation. The situation is explained succinctly in [7, p. 293] as follows (emphasis ours).
The terminology “Suszko’s thesis” has been put forward (see e.g. [26]). This is a quite ambiguous terminology. First all there is [is] a result according to which every logic is bivalent. Such a result is not a thesis. What can be considered as a thesis is rather the interpretation of this result.
Interpretations differ. For example, the ‘real’ reason behind the existence of adequate bivalent semantics for every Tarski-type logical structure - in other words, the ‘logical 2-valuedness’ of Tarski-type logical structures - has been explained by Malinowski in [27, p. 80] as follows.
[L]ogical two-valuedness $\ldots$ is obviously related to the division of the universe of interpretation into two subsets of elements: distinguished and others.
Thus, according to him, ‘logical value’ corresponds to the number of divisions of the universe of interpretation (cf. [38] and [40]). This particular conceptualisation of ‘logical value’ is termed by Malinowski as inferential value. This notion of many-valuedness is different from the more traditional conception of many-valuedness - algebraic many-valuedness [^2] Sometimes ‘algebraic many-valuedness’ is also referred to as ‘referential many-valuedness’ (see, e.g., [28, 29]). However, we will not use this terminology in this chapter.. The difference between these is explained in [37, p. 378] as follows (emphasis ours).
[T]he logical valuations are morphisms (of formulas to the zero-one model) in some exceptional cases, only. Thus, the logical valuations and the algebraic valuations are functions of quite different conceptual nature. The former relate to the truth and falsity, and the latter represent the reference assignments.
Introducing the notion of $q$-consequence operators in [24], Malinowski showed that the logical structures induced by them are inferentially 3-valued, i.e., $\mathbf{ST}$ is false in general. Going further, in [30, 31], he investigates the problem of constructing inferentially $n$-valued logical structures (where $n$ is a natural number) and proposes a partial solution.
Even though much research has been done on inferential many-valuedness, surprisingly there is no explicit or formal definition of the same! [^3] In [35, p. 151], the author defines and examines inferential many-valued inference systems. However, this is not the same as Malinowski’s notion of inferential many-valuedness, as is evident from the following quote. We define and investigate the notion of an inferentially many-valued inference system. In this definition we try to capture the idea of the proof-theoretical reprsentation of truth-functional many-valuedness. On the other hand, the notion of referentially $k$-valued logic, as investigated in the article, can be seen as a special case of our notion algebraically $k$-valued logic (see Definition 2.9). The underlying intuition has been provided several times (see [25, Section 4], [30, Section 2] and [31, Section 3, Section 5] e.g.), of course. The lack of a formal definition, however, is problematic if, e.g., one wants to characterise the class of inferentially $n$-valued structures for a natural number $n$.
One aim of this chapter is to remedy this situation. Based on a rather general notion of semantics (Definition 2.1), we define the notion of inferentially many-valued logical structures (Definition 2.13) and prove that the class of monotonic logical structures can be split into four sub-classes (not necessarily disjoint), each of which has a certain type of adequate $3$-valued semantics (see Theorem 3.8, Theorem 3.15 and Theorem 3.31 for details). Moreover, we identify several classes of logical structures, for which these $3$-valued semantics are respectively minimal (see Theorem 3.10, Theorem 3.17 and Theorem 3.33 for details). It turns out that Malinowski’s notion of many-valuedness is just a particular one among a galaxy of such possible notions (see Remark 2.19). These are the contents of §2 and §3.
§4 is motivated by the observation that most of the research concerning $\mathbf{ST}$ is focused on trying to show that $\mathbf{ST}$ does not hold in general. This has been done either by changing the notion of semantics (e.g. in [24]) or by changing the notion of logical structure (e.g. in [40]). In §4, we introduce a different notion of semantics than the one already mentioned (Definition 4.1) and show that several logical structures can have adequate bivalent semantics. We illustrate this first by providing adequate bivalent semantics for monotonic logical structures and then doing the same for certain particular types of non-monotonic logical structures (Theorem 4.21,Theorem 4.22). We also discuss some problems regarding the existence of adequate bivalent semantics for arbitrary non-monotonic logical structures. Thus, the contents of this section lead to new questions regarding the true nature of logical value, many-valuedness, and most importantly, semantics.
Even though we do not explicitly define logical value, in §5, we define what we mean by a $\kappa$-valued logical structure of order $\lambda$ for any cardinal $\kappa$ and ordinal $\lambda$ (Definition 5.2). For this, we delve into a deeper analysis of the principle of bivalence and explicate how $\mathbf{ST}$ is related to it. The ensuing analysis culminates in the conclusion that many-valued logical structures can be obtained if the bivalence of the higher-order metalogics of the logic under consideration up to a certain level is discarded. This helps to build formal bridges between the theory of graded consequence (see [12]) and the theory of many-valued logical structures, leading to a generalisation of $\mathbf{ST}$.
In the final section, we point out some directions for future research.
2 Fundamental Notions
In the discussions on $\mathbf{ST}$, one important thing that comes out regularly but is not often emphasised may be stated as follows: the justification (or refutation) of $\mathbf{ST}$ depends on the notion of semantics. Usually, the term ‘semantics’ is accompanied by an adjective, e.g., truth-functional semantics, matrix semantics, referential semantics, etc. However, these terms are rather ambiguous because they make sense only if one has a definition of semantics. In [15, p. 384], e.g., Font says the following (emphasis ours).
[…]
So, we propose the following definition of semantics.
Definition 2.1 (Semantics).
Given a set $\mathscr{L}$, a semantics for $\mathscr{L}$ is a tuple $(\mathbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))$, where
  1. (i)
    $\mathbf{M}$ is a set,
  2. (ii)
    $I\neq\emptyset$ and $\models_{i}\,\subseteq\,\mathbf{M}\times\mathcal{P}(\mathscr{L})$ for each $i\in I$,
  3. (iii)
    $\emptyset\subsetneq S\subseteq\{(\models_{i},\models_{j}):(i,j)\in I\times I\}$.
If, in particular, $I=\{1,\ldots,n\}$ with $n\in\mathbf{N}\setminus\{0\}$, instead of writing $(\mathbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))$ we write $(\mathbf{M},\models_{1},\ldots,\models_{n},S,\mathcal{P}(\mathscr{L}))$. If $(m,\Gamma)\in\,\models_{i}$, we write $m\models_{i}\Gamma$. Given a semantics $\mathfrak{S}=(\mathbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))$ the entailment relation induced by $\mathfrak{S}$, $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}\,\subseteq\mathcal{P}(\mathscr{L})\times\mathscr{L}$ is defined as follows.
$\Gamma\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}\alpha\iff\text{for all}~(\models_{i},\models_{j})\in S~\text{and}~m\in\mathbf{M},~m\models_{i}\Gamma~\text{implies that}~m\models_{j}\{\alpha\}$
[…]
Remark 2.2.
Let $(\mathscr{L},W_{\mathfrak{S}})$ be a logical structure induced by the entailment relation corresponding to the semantics $\mathfrak{S}=(\mathbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))$ for $\mathscr{L}$ such that $\mathbf{M}=\emptyset$. If there exists $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$ such that $\alpha\notin W_{\mathfrak{S}}(\Gamma)$, then there exists $m\in\mathbf{M}$ such that $m\models_{i}\Gamma$ but $m\models_{j}\{\alpha\}$. However, since $\mathbf{M}=\emptyset$, $W_{\mathfrak{S}}(\Gamma)=\mathscr{L}$ for all $\Gamma\subseteq\mathscr{L}$.
Example 2.3.
Let $\mathscr{L}$ be a set and $\emptyset\subsetneq\mathcal{V}\subseteq\{0,1\}^{\mathscr{L}}$. Consider the following semantics for $\mathscr{L}$, $\mathfrak{B}_{\mathcal{V}}=(\mathcal{V},\models,S,\mathcal{P}(\mathscr{L}))$, where
  1. (i)
    $\emptyset\subsetneq\mathcal{V}\subseteq\{0,1\}^{\mathscr{L}}$,
  2. (ii)
    $\models\,\,\subseteq\,\mathcal{V}\times\mathcal{P}(\mathscr{L})$,
  3. (iii)
    for all $v\in\mathcal{V}$, $\Gamma\subseteq\mathscr{L}$, $v\models\Gamma$ iff $v(\Gamma)\subseteq\{1\}$,
  4. (iv)
    $S=\{(\models,\models)\}$.
Then, one can show that $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{B}_{\mathcal{V}}})$ is of Tarski-type. In fact, given a Tarski-type logical structure $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$, one can prove that $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,=\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{B}_{\mathcal{S}}}$, where $\mathcal{S}=\{\chi_{W^{\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}}(\Sigma)}\in\{0,1\}^{\mathscr{L}}:\Sigma\subseteq\mathscr{L}\}$.
Example 2.4.
Let $\mathscr{L}$ be a set, $\emptyset\subsetneq\mathcal{V}\subseteq\{0,1\}^{\mathscr{L}}$, and $(P,\preceq)$ is a complete lattice (i.e., a poset where every subset has a supremum as well as an infimum). Consider the following semantics for $\mathscr{L}$, $\mathfrak{P}_{\mathcal{V}}=(\mathscr{L},\{\models\}_{v\in\mathcal{V}},S,\mathcal{P}(\mathscr{L}))$, where
  1. (i)
    for all $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$, $v\in\mathcal{V}$, $\alpha\models_{v}\Gamma$ iff $\inf(v(\Gamma))\leq v(\alpha)$ (here $\inf(v(\Gamma))$ denotes the infimum of $v(\Gamma)$),
  2. (ii)
    $S=\{(\models_{v},\models_{v}):v\in\mathcal{V}\}$.
Then, one can easily show that $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{P}_{\mathcal{V}}})$ is of Tarski-type. In fact, given a Tarski-type logical structure $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$, one can prove that $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,=\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{P}_{\mathcal{S}}}$ where $P=\{0,1\},\preceq\,=\{(0,0),(0,1),(1,1)\}$, and $\mathcal{S}$ is as defined in the previous example.
Remark 2.5.
Given a set $\mathscr{L}$, and a semantics $\mathfrak{S}$ for $\mathscr{L}$, the entailment relation $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}$ as we have defined above, is not the only possible one. Below we mention two other possibilities.
$\Gamma\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}\alpha\iff\text{for all}~(\models_{i},\models_{j})\in S~\text{and}~(m,n)\in\mathbf{M},~m\models_{i}\Gamma~\text{implies that}~n\models_{j}\{\alpha\}$
or,
$\Gamma\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}\alpha\iff\text{there exits}~(\models_{i},\models_{j})\in S~\text{such that for some}~m\in\mathbf{M},~m\models_{i}\Gamma~\text{implies that}~m\models_{j}\{\alpha\}$
However, every logical structure is adequate with respect to some semantics, as the following theorem shows.
Theorem 2.6 (Canonical Semantics of a Logical Structure).
Let $(\mathscr{L},W)$ be a logical structure. Consider the following semantics for $\mathscr{L}$, $\mathfrak{S}=(\mathcal{P}(\mathscr{L}),\models_{1},\models_{2},S,\mathcal{P}(\mathscr{L}))$ where $S=\{(\models_{1},\models_{2})\}$, and for all $\Gamma\cup\Sigma\subseteq\mathscr{L}$ $\models_{1}$ and $\models_{2}$ are defined respectively as follows:
$\Sigma\models_{1}\Gamma~\text{iff}~\Sigma=\Gamma$
and
$\Sigma\models_{2}\Gamma~\text{iff}~\Gamma\subseteq W(\Sigma).$
Then, $W=W_{\mathfrak{S}}$.
Proof.
Suppose the contrary, i.e., suppose that $W\neq W_{\mathfrak{S}}$. Then, for some $\Gamma\subseteq\mathscr{L}$, $W(\Gamma)\neq W_{\mathfrak{S}}(\Gamma))$. If possible, let $W(\Gamma)\not\subseteq W_{\mathfrak{S}}(\Gamma)$. Then there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W_{\mathfrak{S}}(\Gamma)$. So, there exists $\Sigma\subseteq\mathscr{L}$ such that $\Sigma\models_{1}\Gamma$ but $\Sigma\not\models_{2}\{\alpha\}$. Now, $\Sigma\models_{1}\Gamma$ implies that $\Sigma=\Gamma$. But then, since $\Sigma\not\models_{2}\{\alpha\}$, we must have $\{\alpha\}\not\subseteq W(\Gamma)$, i.e., $\alpha\notin W(\Gamma)$ - a contradiction. Thus, we $W_{\mathfrak{S}}(\Gamma)\not\subseteq W(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$.
So, there exists $\alpha\in W_{\mathfrak{S}}(\Gamma)$ such that $\alpha\notin W(\Gamma)$. Choose this $\Gamma$, and note that $\Gamma\models_{1}\Gamma$. Therefore, since $\alpha\in W_{\mathfrak{S}}(\Gamma)$, we must have $\Gamma\models_{2}\{\alpha\}$, i.e., $\alpha\in W(\Gamma)$ - a contradiction. Thus, $W_{\mathfrak{S}}(\Gamma)=W(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$, i.e. $W=W_{\mathfrak{S}}$. ∎
Remark 2.7.
Generally, the notion of semantics is not defined explicitly. A notable exception [^4] Equivalent notions under the names of model-based structure and abstract model logic have been defined in [39, p. 106] and [21, p. 21] respectively. is [4, pp. 107-108] where a semantics is defined as a triple $\langle\mathfrak{F};\boldsymbol{U};\text{mod}\rangle$, where
  • $\mathfrak{F}$ is a set, called the domain of the semantics,
  • $\boldsymbol{U}$ is a set, called the universe of the semantics,
  • mod is a map from $\mathfrak{F}$ to $\mathcal{P}(\boldsymbol{U})$:
    $\text{mod}:\mathfrak{F}$ $\to\mathcal{P}(\boldsymbol{U}),$
    $F$ $\mapsto\text{mod}(F)$
which can be extended to a function $\textbf{mod}:\mathcal{P}(\mathfrak{F})\to\mathcal{P}(\boldsymbol{U})$ as follows:
$\bf{mod}:\mathcal{P}(\mathfrak{F})$ $\to\mathcal{P}(\boldsymbol{U}),$
$T$ $\mapsto{\bf{mod}}(T)=\displaystyle\bigcap_{F\in T}\text{mod}(F)$
Any semantics $\langle\mathfrak{F};\boldsymbol{U};\text{mod}\rangle$ induces a semantics (in the sense of Definition 2.1) $(\mathbf{M},\models,S,\mathcal{P}(\mathfrak{F}))$ where $\mathbf{M}=\mathcal{P}(\boldsymbol{U})$, $S=\{(\models,\models)\}$ and for all $\Sigma\subseteq\mathfrak{F}$, $\Gamma\subseteq\boldsymbol{U}$, $\Gamma\models\Sigma$ iff $\mathbf{mod}(\Sigma)=\Gamma$.
By Theorem 2.6, every logical structure has an adequate semantics. Even though this result might be of some theoretical interest, the construction of the semantics may seem ad hoc. We, therefore, introduce several less abstract notions of semantics and investigate their inter-relations in some detail. We will return to a deeper analysis of results like this in §5 when we discuss the relationships between different formulations of the principle of bivalence and $\mathbf{ST}$.
Definition 2.8 (Algebraically $\kappa$-valued Semantics).
Given an algebra $\mathbscr{L}$, a semantics $\mathfrak{S}(=(\mathbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathbscr{L})))$ for $\mathbscr{L}$, and a cardinal $\kappa$, $\mathfrak{S}$ is said to be algebraically $\kappa$-valued if there exists an algebra $\mathbscr{A}$ of the same similarity type as $\mathbscr{L}$, with universe $A$ such that the following properties hold.
  1. (i)
    $\mathbf{M}$ is the set of all homomorphisms from $\mathbscr{L}$ to $\mathbscr{A}$.
  2. (ii)
    $|A|=\kappa$.
  3. (iii)
    $\bigcup_{m\in\mathbf{M}}m(\mathscr{L})\subseteq A$.
  4. (iv)
    For all $m\in\mathbf{M},i\in I$ and $\Gamma\subseteq\mathscr{L}$,
    $m\models_{i}\Gamma~\text{iff}~m(\Gamma)\subseteq D_{i}$
    where $\emptyset\subsetneq D_{i}$ for each $i\in I$ and $\bigcup_{i\in I}D_{i}\subsetneq A$.
Given an algebra $\mathbscr{L}$, the class of all algebraically $\kappa$-valued semantics for $\mathbscr{L}$ will be denoted by $\mathcal{A}_{\kappa}(\mathbscr{L})$.
Definition 2.9 (Algebraically $\kappa$-valued Logic).
Given a cardinal $\kappa$, a logic $(\mathbscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$ (i.e., where $\mathbscr{L}$ is an algebra) is said to be algebraically $\kappa$-valued if there exists an $\mathfrak{S}=(\mathbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))\in\mathcal{A}_{\kappa}(\mathbscr{L})$ such that the following statements hold.
  1. (i)
[…]
  2. (ii)
    $\kappa>1$
  3. (iii)
[…]
A class of logics $\mathcal{K}$ is said to be algebraically $\kappa$-valued if it contains an algebraically $\kappa$-valued logic.
Remark 2.10.
In other words, a logic $(\mathbscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$ is said to be algebraically $\kappa$-valued if it is adequate with respect to an algebraically $\kappa$-valued semantics, where
This can be seen as a generalisation of the usual understanding of $\kappa$-valued logics (see, e.g., [5] and [32]).
Definition 2.11 (Functionally $\kappa$-valued Semantics).
Given a set $\mathscr{L}$, a semantics $\mathfrak{S}=(\mathbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))$ for $\mathscr{L}$, and a cardinal $\kappa$, $\mathfrak{S}$ is said to be functionally $\kappa$-valued if it satisfies the following properties.
  1. (i)
    $\emptyset\subsetneq\mathbf{M}\subseteq A^{\mathscr{L}}$ for some nonempty set $A$.
  2. (ii)
    $|A|=\kappa$.
  3. (iii)
    $\bigcup_{m\in\mathbf{M}}m(\mathscr{L})\subseteq A$.
  4. (iv)
    For all $m\in\mathbf{M},i\in I$ and $\Gamma\subseteq\mathscr{L}$,
    $m\models_{i}\Gamma~\text{iff}~m(\Gamma)\subseteq D_{i}$
    where $\emptyset\subsetneq D_{i}$ for each $i\in I$ and $\bigcup_{i\in I}D_{i}\subsetneq A$.
Given a set $\mathscr{L}$, the class of all functionally $\kappa$-valued semantics for $\mathscr{L}$ will be denoted by $\mathcal{F}_{\kappa}(\mathscr{L})$.
Remark 2.12.
The entailment relation induced by a functionally $\kappa$-valued semantics (for a given cardinal $\kappa$) is more general than the ones induced by $t$-, $f$-, $p$- and $q$-entailment relations (see [9, p. 238] for details).
Definition 2.13 (Inferentially $\kappa$-valued Logical Structure).
Given a cardinal $\kappa$, a logical structure $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$ said to be inferentially $\kappa$-valued if there exists $\mathfrak{S}(=(\mathbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L})))\in\mathcal{F}_{\kappa}(\mathscr{L})$ such that the following statements hold.
  1. (i)
[…]
  2. (ii)
    $\kappa>1$
  3. (iii)
[…]
A class of logical structures $\mathcal{K}$ is said to be inferentially $\kappa$-valued if it contains an inferentially $\kappa$-valued logical structure.
Definition 2.14 (Granular Semantics).
A semantics $\mathfrak{S}=(\mathbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))$ for $\mathscr{L}$ is said to be granular at the $i$-th component if for all $\Gamma\subseteq\mathscr{L},m\in\mathbf{M}$, the following statement holds.
$m\models_{i}\Gamma~\text{iff}~m\models_{i}\{\alpha\}~\text{for all}~\alpha\in\Gamma$
If $\mathfrak{S}$ is granular at the $i$-th component for all $i\in I$, then $\mathfrak{S}$ is said to be strongly granular for $\mathscr{L}$.
Remark 2.15.
For all cardinal $\kappa$, every functionally $\kappa$-valued semantics is strongly granular. The converse of this statement forms the basis of a general form of Suszko Reduction, which we discuss in §4.
Remark 2.16.
We would like to point out that the definition of inferentially $\kappa$-valued logical structures, as we have given in Definition 2.13 is not the same as Malinowski’s treatment of the same. Nevertheless, we have chosen to adopt the same terminology because we think that this definition captures the essential idea: a logical structure $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$ is inferentially $\kappa$-valued if it is adequate with respect to a functionally $\kappa$-valued semantics, where
Remark 2.17.
Notice that both Definition 2.9 and Definition 2.13 can be understood independent of the precise definition of $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}$. Hence, even though, e.g., Definition 2.13 implies that the every inferentially $\kappa$-valued logical structure is necessarily monotonic  [^5] A special thanks to Laurent Dion for pointing this out. (see Theorem 3.1 for the proof) – hence, inferentially 4-valued, by Corollary 3.41 – it is possible to formulate other definitions of ‘inferentially $\kappa$-valued logical structure’ (depending on different definitions of $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{\mathfrak{S}}$), that are not necessarily so. Similar remarks apply to Definition 2.9.
Remark 2.18.
Let $(\mathbscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$ be a logic that is algebraically $\kappa$-valued. Suppose that the corresponding logical structure $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$ is inferentially $\lambda$-valued, and is adequate with respect to an functionally $\kappa$-valued semantics. Consequently, by the previous remark, it follows that $\lambda\leq\kappa$. In other words, in general, the logical structure corresponding to algebraically $\kappa$-valued logic is inferentially $\lambda$-valued for $\lambda\leq\kappa$.
Remark 2.19.
It is possible to introduce more general notions of ‘many-valuedness’ by considering objects more general than functions in (i) of Definition 2.11. For example, one may consider the set of all relations, instead of the set of all functions and obtain the corresponding notion of many-valuedness. The upshot of this discussion is that both the standard as well as Malinowski’s notion of inferential many-valuedness are just two particular conceptions of many possible ones. However, since in this chapter, we will be focusing only on inferential many-valuedness, we don’t pursue this issue in more detail.
3 Monotonic Logical Structures: A Case Study
In this section, we consider monotonic logical structures. We show that the class of such logical structures can be split into four subclasses, each functionally 3-valued at least. Later, in §§4.1, we show how each can be given a bivalent characterisation.
Interestingly enough, the entailment relations induced by strongly granular semantics give rise to monotonic logical structures. This is the content of the following theorem.
Theorem 3.1 (Representation Theorem for Monotonic Logical Structures (Part I)).
Let $\mathfrak{S}=(\textbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))$, be a strongly granular semantics for $\mathscr{L}$. Then $W_{\mathfrak{S}}$ is monotonic.
Proof.
If possible, suppose $\Gamma\cup\Sigma\subseteq\mathscr{L}$ be such that $\Gamma\subseteq\Sigma$ but $W_{\mathfrak{S}}(\Gamma)\not\subseteq W_{\mathfrak{S}}(\Sigma)$. Hence, there exists $\alpha\in W_{\mathfrak{S}}(\Gamma)$ such that $\alpha\notin W_{\mathfrak{S}}(\Sigma)$. Since $\alpha\notin W_{\mathfrak{S}}(\Sigma)$, there exists $(i,j)\in I\times I$ with $(\models_{i},\models_{j})\in S$ and $m\in\mathbf{M}$ such that $m\models_{i}\Sigma$ but $m\not\models_{j}\{\alpha\}$. However, since $\Gamma\subseteq\Sigma$, $m\models_{i}\Sigma$ and $\mathfrak{S}$ is strongly granular, it follows that $m\models_{i}\Gamma$ as well. But then, since $\alpha\in W_{\mathfrak{S}}(\Gamma)$ and $(\models_{i},\models_{j})\in S$, we must have $m\models_{j}\{\alpha\}$ - a contradiction. Thus, $W_{\mathfrak{S}}(\Gamma)\subseteq W_{\mathfrak{S}}(\Sigma)$. Hence, $W_{\mathfrak{S}}$ is monotonic. ∎
Every logical structure induced by a monotonic consequence operator has a functionally 4-valued characterisation, as we show below. [^6] An equivalent version of the same result, in a different setting, can be obtained by combining [9, Theorem 2 and 11].
Theorem 3.2 (Representation Theorem for Monotonic Logical Structures (Part II)).
Let $(\mathscr{L},W)$ be a monotonic logical structure and $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S,\mathcal{P}(\mathscr{L}))$ is a semantics for $\mathscr{L}$ satisfying the following properties.
  1. $\bullet$
    $\mathbf{M}=\{\mu_{\Sigma}\in\{0,1,2,3\}^{\mathscr{L}}:\Sigma\subseteq\mathscr{L}\}$ where
    $\mu_{\Sigma}(\beta)=\begin{cases}0&\text{if}~\beta\notin W(\Sigma)\cup\Sigma\\ 1&\text{if}~\beta\in\Sigma\cap W(\Sigma)\\ 2&\text{if}~\beta\in\Sigma\setminus W(\Sigma)\\ 3&\text{if}~\beta\in W(\Sigma)\setminus\Sigma\end{cases}$
  2. $\bullet$
    For all $m\in\mathbf{M},\Gamma\subseteq\mathscr{L}$,
    $m\models_{1}\Gamma~\text{iff}~m(\Gamma)\subseteq\{1,2\}$
    $m\models_{2}\Gamma~\text{iff}~m(\Gamma)\subseteq\{1,3\}$
  3. $\bullet$
    $S=\{(\models_{1},\models_{2})\}$.
Then, $W=W_{\mathfrak{S}}$.
Proof.
Suppose the contrary, i.e., suppose that $W\neq W_{\mathfrak{S}}$. Then there exists $\Gamma\subseteq\mathscr{L}$ such that $W(\Gamma)\neq W_{\mathfrak{S}}(\Gamma)$. Therefore, $W(\Gamma)\not\subseteq W_{\mathfrak{S}}(\Gamma)$ or $W_{\mathfrak{S}}(\Gamma)\not\subseteq W(\Gamma)$.
If $W(\Gamma)\not\subseteq W_{\mathfrak{S}}(\Gamma)$, then there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W_{\mathfrak{S}}(\Gamma)$. So, there exists $\mu_{\Sigma}\in\mathbf{M}$ such that $\mu_{\Sigma}\models_{1}\Gamma$ but $\mu_{\Sigma}\not\models_{2}\{\alpha\}$. Now, $\mu_{\Sigma}\models_{1}\Gamma$ implies that $\mu_{\Sigma}(\Gamma)\subseteq\{1,2\}$, or, equivalently $\Gamma\subseteq\Sigma$. Then, since $W$ is monotonic, we have $W(\Gamma)\subseteq W(\Sigma)$. Consequently, $\alpha\in W(\Gamma)$ implies that $\alpha\in W(\Sigma)$. But, $\alpha\in W(\Sigma)$ is equivalent to requiring that $\mu_{\Sigma}(\alpha)\in\{1,3\}$, i.e., $\mu_{\Sigma}\models_{2}\{\alpha\}$ - a contradiction. Consequently, $W(\Gamma)\subseteq W_{\mathfrak{S}}(\Gamma)$.
If, on the other hand, $W_{\mathfrak{S}}(\Gamma)\not\subseteq W(\Gamma)$, then there exists $\alpha\in W_{\mathfrak{S}}(\Gamma)$ such that $\alpha\notin W(\Gamma)$. Now, $\alpha\notin W(\Gamma)$ is equivalent to saying that $\mu_{\Gamma}\not\models_{2}\{\alpha\}$. However, since $\mu_{\Gamma}\models_{1}\Gamma$ and $\alpha\in W_{\mathfrak{S}}(\Gamma)$, we must have $\mu_{\Gamma}\models_{2}\{\alpha\}$ - a contradiction. Consequently, $W_{\mathfrak{S}}(\Gamma)\subseteq W(\Gamma)$ as well. This, however, is contrary to our assumption that $W\neq W_{\mathfrak{S}}$. Thus, $W=W_{\mathfrak{S}}$. ∎
Remark 3.3.
The above result can be rephrased as saying that every monotonic logical structure has an adequate functionally $4$-valued semantics. This can be seen by considering $A=\{0,1,2,3\}$, $\kappa=4$, $D_{1}=\{1,2\}$ and $D_{2}=\{1,3\}$ in Definition 2.11.
We now show that every monotonic logical structure is induced by a $q$-, $p$-, $s_{\kappa}$- or $r_{\kappa}$-consequence operator (the last two consequence operators will be defined respectively in §§3.3 and §§3.3).
3.1 $q$-consequence Operators
As mentioned earlier, $q$-consequence operators were introduced by Malinowski to show that $\mathbf{ST}$ does not apply, in general, to the logical structures induced by these operators, provided ‘logical $n$-valuedness’ is understood as ‘inferential $n$-valuedness’. In this section, we delve into a detailed study of $q$-consequence operators and show that every logical structure induced by $q$-consequence operator has an adequate functionally $3$-valued semantics.
Definition 3.4 ($q$-consequence Operator).
A $q$-consequence operator on $\mathscr{L}$ is a function $W:\mathcal{P}(\mathscr{L})\to\mathcal{P}(\mathscr{L})$ satisfying the following properties.
  1. (i)
    For all $\Gamma\cup\Sigma\subseteq\mathscr{L}$, if $\Gamma\subseteq\Sigma$ then $W(\Gamma)\subseteq W(\Sigma)$. (Monotonicity)
  2. (ii)
    For all $\Gamma\subseteq\mathscr{L}$ $W(W(\Gamma)\cup\Gamma)=W(\Gamma)$. (Quasi-closure)
A logical structure $(\mathscr{L},W)$ is said to be of q-type if $W$ is a $q$-consequence operator on $\mathscr{L}$.
We now prove a characterisation theorem for $q$-type logical structures using downward $q$-closed sets, defined below. One might compare this result with [34, Theorem 2.8], and downward $q$-closed sets with strongly closed sets [34, Definition 2.6].
Definition 3.5 (Downward $q$-closed Set).
Let $(\mathscr{L},W)$ be a logical structure and $\Delta\subseteq\mathscr{L}$. The set $\Delta$ is said to be downward $q$-closed if for all $\Sigma\subseteq W^{\infty}(\Delta)$, $W(\Sigma)\subseteq W(\Delta)$. Here $W^{0}(\Delta)=\Delta$ and $W^{\infty}(\Delta)=\displaystyle\bigcup_{i=0}^{\infty}W^{i}(\Delta)$.
Theorem 3.6 (Characterisation of $q$-type).
Let $(\mathscr{L},W)$ be a logical structure. Then, the following statements are equivalent.
  1. (1)
    $W$ is a $q$-consequence operator.
  2. (2)
    $W$ is monotonic, and for all $\Gamma\subseteq\mathscr{L}$, $W\left(W^{\infty}(\Gamma)\right)=W(\Gamma)$. (For all $\Gamma\subseteq\mathscr{L}$, $W^{0}(\Gamma)=\Gamma$.)
  3. (3)
    For every $\Delta\subseteq\mathscr{L}$, $\Delta$ is downward $q$-closed.
  4. (4)
    For all $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$, if $\alpha\notin W(\Gamma)$ then there exists a downward $q$-closed $\Sigma\supseteq\Gamma$ such that $\alpha\notin W(\Sigma)$.
  • (5)
    For all $\Gamma\cup\Sigma\subseteq\mathscr{L}$, $\Gamma\subseteq W(\Sigma)\cup\Sigma$ implies that $W(\Gamma)\subseteq W(\Sigma)$.
    Proof.
    $\underline{(1)\implies(2):}$ Since $W$ is a $q$-consequence operator, it is monotonic. So, it suffices to show that $W\left(W^{\infty}(\Gamma)\right)=W(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$. Let $\Gamma\subseteq\mathscr{L}$. Then,
    $W^{2}(\Gamma)$ $=W(W(\Gamma))$
    $\subseteq W(W(\Gamma)\cup\Gamma)$ $\text{(since}~W~\text{is monotonic)}$
    $=W(\Gamma)$ $\text{(since}~W~\text{satisfies quasi-closure)}$
    Moreover, if for some $i\geq 1$ $W^{i}(\Gamma)\subseteq W(\Gamma)$ then $W^{i+1}(\Gamma)=W(W^{i}(\Gamma))\subseteq W(W(\Gamma))\subseteq W(\Gamma)$. Thus, for all $i\geq 1$, $W^{i}(\Gamma)\subseteq W(\Gamma)$. This implies $\bigcup_{i=1}^{\infty}W^{i}(\Gamma)\subseteq W(\Gamma)$, and hence
[…]
    $\underline{(2)\implies(1):}$ Since $W$ is monotonic, it suffices to show that $W$ satisfies quasi-closure. For this, let $\Gamma\subseteq\mathscr{L}$ and consider $W(W(\Gamma)\cup\Gamma)$. Notice that $\Gamma\subseteq W(\Gamma)\cup\Gamma$ and $W(\Gamma)\cup\Gamma\subseteq W^{\infty}(\Gamma)$. Consequently,
[…]
    So, $W$ satisfies quasi-closure, and hence is a $q$-consequence operator. Therefore, (1) holds.
    $\underline{(2)\implies(3):}$ Let $\Delta\subseteq\mathscr{L}$ and $\Sigma\subseteq W^{\infty}(\Delta)$. Then, by monotonicity of $W$ and by (2), we have $W(\Sigma)\subseteq W\left(W^{\infty}(\Delta)\right)=W(\Delta)$. Thus, (3) holds.
    $\underline{(3)\implies(2):}$ Suppose that $W$ is not monotonic. So, there exists $\Gamma\cup\Sigma\subseteq\mathscr{L}$ with $\Gamma\subseteq\Sigma$ such that $W(\Gamma)\not\subseteq W(\Sigma)$. This implies that there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W(\Sigma)$. Since $\Gamma\subseteq\Sigma\subseteq W^{\infty}(\Sigma)$ and $\Sigma$ itself is downward $q$-closed by (3), it follows that $W(\Gamma)\subseteq W(\Sigma)$. Hence, $\alpha\in W(\Gamma)$ implies that $\alpha\in W(\Sigma)$ as well - a contradiction. Therefore, $W$ is monotonic.
    Now, let $\Gamma\subseteq\mathscr{L}$. Since $\Gamma\subseteq W^{\infty}(\Gamma)$, $W(\Gamma)\subseteq W\left(W^{\infty}(\Gamma)\right)$ by monotonicity of $W$. Since $\Gamma$ is downward $q$-closed by (3) and $W^{\infty}(\Gamma)\subseteq W^{\infty}(\Gamma)$, trivially, $W\left(W^{\infty}(\Gamma)\right)\subseteq W(\Gamma)$. Hence (2) holds.
    $\underline{(2)\implies(4):}$ Let $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$ be such that $\alpha\notin W(\Gamma)$. Clearly $\Gamma\subseteq W^{\infty}(\Gamma)$. By (2), $W(W^{\infty}(\Gamma))=W(\Gamma)$. So, $\alpha\notin W(W^{\infty}(\Gamma))$. Thus, it suffices to show that $W^{\infty}(\Gamma)$ is downward $q$-closed.
    Now, it follows that for all $i\geq 1$ we have, $W^{i}\left(W^{\infty}(\Gamma)\right)=W^{i}(\Gamma)$. Hence, $W^{\infty}\left(W^{\infty}(\Gamma)\right)=W^{\infty}(\Gamma)$.
    Let $\Lambda\subseteq W^{\infty}\left(W^{\infty}(\Gamma)\right)$. Then $\Lambda\subseteq W^{\infty}(\Gamma)$. Now, since $W$ is monotonic, $W(\Lambda)\subseteq W\left(W^{\infty}(\Gamma)\right)$ for all $\Lambda\subseteq W^{\infty}(\Gamma)$. This implies that $W^{\infty}(\Gamma)$ is downward $q$-closed. This proves (4).
    $\underline{(4)\implies(2):}$ We first show that $W$ is monotonic. For this, suppose that there exists $\Gamma\cup\Sigma\subseteq\mathscr{L}$ with $\Gamma\subseteq\Sigma$ such that $W(\Gamma)\not\subseteq W(\Sigma)$. This implies that there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W(\Sigma)$. Consequently, by (4), there exists $\Delta\supseteq\Sigma$ such that $\Delta$ is downward $q$-closed and $\alpha\notin W(\Delta)$. Now, since $\Gamma\subseteq\Sigma\subseteq\Delta\subseteq W^{\infty}(\Delta)$, and $\Delta$ is downward $q$-closed, $W(\Gamma)\subseteq W(\Delta)$. So, since $\alpha\in W(\Gamma)$, $\alpha\in W(\Delta)$ as well - a contradiction. Hence, $W$ is monotonic.
    We now show that $W$ satisfies quasi-closure as well. Suppose the contrary, i.e., $W(W^{\infty}(\Gamma))\neq W(\Gamma)$. Then we must have $W\left(W^{\infty}(\Gamma)\right)\not\subseteq W(\Gamma)$ for some $\Gamma\subseteq\mathscr{L}$. So, there exists $\alpha\in W\left(W^{\infty}(\Gamma)\right)$ such that $\alpha\notin W(\Gamma)$. Then, by (4), there exists a downward $q$-closed $\Sigma\supseteq\Gamma$ such that $\alpha\notin W(\Sigma)$. Since $\Sigma$ is downward $q$-closed and $W^{\infty}(\Sigma)\subseteq W^{\infty}(\Sigma)$, it follows that $\alpha\notin W\left(W^{\infty}(\Sigma)\right)$. Now, as $W$ is monotonic, $\Gamma\subseteq\Sigma$ implies that $W^{\infty}(\Gamma)\subseteq W^{\infty}(\Sigma)$, and hence $W\left(W^{\infty}(\Gamma)\right)\subseteq W\left(W^{\infty}(\Sigma)\right)$. Thus, $\alpha\notin W\left(W^{\infty}(\Gamma)\right)$ as well - a contradiction. Hence, $W\left(W^{\infty}(\Gamma)\right)\subseteq W(\Gamma)$. This proves (2).
    $\underline{(1)\implies(5):}$ Let $\Gamma\cup\Sigma\subseteq\mathscr{L}$ be such that $\Gamma\subseteq W(\Sigma)\cup\Sigma$. Then,
    $\Gamma\subseteq\displaystyle W(\Sigma)\cup\Sigma$ $\implies W(\Gamma)\subseteq W(W(\Sigma)\cup\Sigma)$ (by monotonicity)
    $\implies W(\Gamma)\subseteq W(\Sigma)$ (by quasi-closure)
    Thus $(5)$ holds.
    $\underline{(5)\implies(1):}$ Let $\Gamma\cup\Sigma\subseteq\mathscr{L}$ be such that $\Gamma\subseteq\Sigma$. Then, $\Gamma\subseteq W(\Sigma)\cup\Sigma$ which implies that $W(\Gamma)\subseteq W(\Sigma)$.
    Hence, $W$ is monotonic. Now, for any $\Gamma\subseteq\mathscr{L}$, $\Gamma\subseteq W(\Gamma)\cup\Gamma$, and hence by monotonicity of $W$, $W(\Gamma)\subseteq W\left(W(\Gamma)\cup\Gamma\right)$. The reverse inclusion is immediate since, by (5), $W(\Gamma)\cup\Gamma\subseteq W(\Gamma)\cup\Gamma$ implies that $W\left(W(\Gamma)\cup\Gamma\right)\subseteq W(\Gamma)$. Hence, $W$ is a $q$-consequence operator. This proves (1). ∎
    Theorem 3.7 (Representation Theorem for $q$-type (Part I)).
    Let $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S_{q},\mathcal{P}(\mathscr{L}))$ be a strongly granular semantics for $\mathscr{L}$ such that $\models_{2}\,\subseteq\,\models_{1}$ and $S_{q}=\{(\models_{1},\models_{2})\}$. Then $W_{\mathfrak{S}}$ is a $q$-consequence operator.
    Proof.
    Monotonicity of $W_{\mathfrak{S}}$ follows from Theorem 3.1. Hence, it suffices to show that $W_{\mathfrak{S}}$ satisfies the quasi-closure. Suppose the contrary, i.e., suppose that for some $\Gamma\subseteq\mathscr{L}$, $W_{\mathfrak{S}}(W_{\mathfrak{S}}(\Gamma)\cup\Gamma)\neq W_{\mathfrak{S}}(\Gamma)$. Now, $W_{\mathfrak{S}}(\Gamma)\subseteq W_{\mathfrak{S}}(W_{\mathfrak{S}}(\Gamma)\cup\Gamma)$ since $W_{\mathfrak{S}}$ is monotonic. Thus, $W_{\mathfrak{S}}(W_{\mathfrak{S}}(\Gamma)\cup\Gamma)\not\subseteq W_{\mathfrak{S}}(\Gamma)$. So, there exists $\alpha\in W_{\mathfrak{S}}(W_{\mathfrak{S}}(\Gamma)\cup\Gamma)$ such that $\alpha\notin W_{\mathfrak{S}}(\Gamma)$. So, there exists $m\in\mathbf{M}$ such that $m\models_{1}\Gamma$ but $m\not\models_{2}\{\alpha\}$.
    Let $\beta\in W_{\mathfrak{S}}(\Gamma)$. Then, since $m\models_{1}\Gamma$, we have $m\models_{2}\{\beta\}$. Furthermore, since $\models_{2}\,\subseteq\,\models_{1}$ and $m\models_{2}\{\beta\}$, it follows that $m\models_{1}\{\beta\}$ as well. Thus, $m\models_{1}\{\beta\}$. Now, since $m\models_{1}\Gamma$ and $\mathfrak{S}$ is strongly granular, $m\models_{1}\{\gamma\}$ for all $\gamma\in\Gamma$. Hence $m\models_{1}\{\sigma\}$ for all $\sigma\in W_{\mathfrak{S}}(\Gamma)\cup\Gamma$. Again, since $\mathfrak{S}$ is strongly granular, $m\models_{1}W_{\mathfrak{S}}(\Gamma)\cup\Gamma$. But then, since $\alpha\in W_{\mathfrak{S}}(W_{\mathfrak{S}}(\Gamma)\cup\Gamma)$, we must have $m\models_{2}\{\alpha\}$ - a contradiction. Thus, $W_{\mathfrak{S}}(W_{\mathfrak{S}}(\Gamma)\cup\Gamma)\subseteq W_{\mathfrak{S}}(\Gamma)$. Hence, $W_{\mathfrak{S}}$ satisfies the quasi-closure property as well, and therefore is a $q$-consequence operator. ∎
    Let $(\mathbscr{L},W)$ be a logic where $\mathbscr{L}$ is an algebra with universe $\mathscr{L}$. It is said to satisfy structurality if for all homomorphisms $\sigma:\mathbscr{L}\to\mathbscr{L}$ and all $\Gamma\subseteq\mathscr{L}$, $\sigma(W(\Gamma))\subseteq W(\sigma(\Gamma))$. If a logic satisfies structurality, we call it structural (see [16, 17] for more details and context).
    The proof that every structural $q$-type logic has a 3-valued description is already been given, e.g., in [30, pp. 204-206]. However, a similar argument works for arbitrary $q$-type logical structures as well. Below we give a proof of the same, to keep the presentation self-contained.
    Theorem 3.8 (Representation Theorem for $q$-type (Part II)).
    Let $(\mathscr{L},W)$ be a $q$-type logical structure. Let $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S_{q},\mathcal{P}(\mathscr{L}))$ be a semantics for $\mathscr{L}$ satisfying the following properties.
    1. $\bullet$
      $\mathbf{M}=\{\mu_{\Sigma}\in\{0,1,2\}^{\mathscr{L}}:\Sigma\subseteq\mathscr{L}\}$, where
      $\mu_{\Sigma}(\beta)=\begin{cases}0&\text{if}~\beta\notin W(\Sigma)\cup\Sigma\\ 1&\text{if}~\beta\in W(\Sigma)\\ 2&\text{if}~\beta\in(W(\Sigma)\cup\Sigma)\setminus W(\Sigma)\end{cases}$
    2. $\bullet$
      For all $m\in\mathbf{M},\Gamma\subseteq\mathscr{L}$,
      $m\models_{1}\Gamma~\text{iff}~m(\Gamma)\subseteq\left\{1,2\right\}$
      $m\models_{2}\Gamma~\text{iff}~m(\Gamma)\subseteq\{1\}$
    3. $\bullet$
      $S_{q}$ is as defined in the previous theorem.
    Then, $W=W_{\mathfrak{S}}$.
    Proof.
    Suppose, $\Gamma\subseteq\mathscr{L}$ is such that $W(\Gamma)\not\subseteq W_{\mathfrak{S}}(\Gamma)$. Then there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W_{\mathfrak{S}}(\Gamma)$. Hence, there exists $m\in\textbf{M}$ such that $m\models_{1}\Gamma$ but $m\not\models_{2}\{\alpha\}$. Now, there exists $\Sigma\subseteq\mathscr{L}$ such that $m=\mu_{\Sigma}$. Then, $\mu_{\Sigma}\models_{1}\Gamma$ implies that $\mu_{\Sigma}(\Gamma)\subseteq\{1,2\}$, or equivalently $\Gamma\subseteq W(\Sigma)\cup\Sigma$. Since $W$ is a $q$-consequence operator, it is monotonic. So, $W(\Gamma)\subseteq W(W(\Sigma)\cup\Sigma)$. Moreover, since $W$ satisfies quasi-closure, it follows that $W(W(\Sigma)\cup\Sigma)=W(\Sigma)$. Consequently, $W(\Gamma)\subseteq W(\Sigma)$. Then, as $\alpha\in W(\Gamma)$, $\alpha\in W(\Sigma)$. But then $\mu_{\Sigma}(\alpha)=1$, i.e., $\mu_{\Sigma}\models_{2}\{\alpha\}$, i.e., $m\models_{2}\{\alpha\}$ - a contradiction. Hence, $W(\Gamma)\subseteq W_{\mathfrak{S}}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$.
    Next, let $\Gamma\subseteq\mathscr{L}$ and $\alpha\in W_{\mathfrak{S}}(\Gamma)$. Now, $\mu_{\Gamma}(\Gamma)\subseteq\{1,2\}$, i.e., $\mu_{\Gamma}\models_{1}\Gamma$. Therefore, as $\alpha\in W_{\mathfrak{S}}(\Gamma)$, it follows that $\mu_{\Gamma}\models_{2}\{\alpha\}$, which implies that $\alpha\in W(\Gamma)$. Thus, $W_{\mathfrak{S}}(\Gamma)\subseteq W(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$. Hence $W=W_{\mathfrak{S}}$. ∎
    Theorem 3.9 (Representation Theorem for $q$-type (Part III)).
    Let $(\mathscr{L},W)$ be a logical structure where $W$ is a $q$-consequence operator. Let $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S,\mathcal{P}(\mathscr{L}))$ where $S=\{(\models_{1},\models_{2})\}$ is a strongly granular semantics for $\mathscr{L}$. If $W=W_{\mathfrak{S}}$, then $\models_{1}$ and $\models_{2}$ satisfy the following property:
    For all $\Gamma\cup\Sigma\cup\{\alpha\}\subseteq\mathscr{L}$ such that $\Gamma\subseteq W(\Sigma)\cup\Sigma$, $\pi_{1}(\models_{1}\cap~(\mathbf{M}\times\{\Sigma\}))\not\subseteq\,\pi_{1}(\models_{2}\cap~(\mathbf{M}\times\{\{\alpha\}\}))$ implies that $\pi_{1}(\models_{1}\cap~(\mathbf{M}\times\{\Gamma\}))\not\subseteq\,\pi_{1}(\models_{2}\cap~(\mathbf{M}\times\{\{\alpha\}\}))$.
    Here $\pi_{1}:\mathbf{M}\times\mathcal{P}(\mathscr{L})\to\mathbf{M}$ is defined as follows: $\pi_{1}(m,\Delta)=m$ for all $(m,\Delta)\in\mathbf{M}\times\mathcal{P}(\mathscr{L})$.
    Proof.
    Let $\Gamma\cup\Sigma\subseteq\mathscr{L}$ be such that $\Gamma\subseteq W(\Gamma)\cup\Gamma$. Therefore, since $W$ is a $q$-consequence operator, by Theorem 3.6(5) we have $W(\Gamma)\subseteq W(\Sigma)$. Since $W=W_{\mathfrak{S}}$, this implies $W_{\mathfrak{S}}(\Gamma)\subseteq W_{\mathfrak{S}}(\Sigma)$.
    Let $\alpha\in\mathscr{L}$ be such that $\pi_{1}(\models_{1}\cap~(\mathbf{M}\times\{\Sigma\}))\not\subseteq\,\pi_{1}(\models_{2}\cap~(\mathbf{M}\times\{\{\alpha\}\}))$. Then, for some $m\in\mathbf{M}$, $m\in\pi_{1}(\models_{1}\cap~(\mathbf{M}\times\{\Sigma\}))$ but $m\notin\pi_{1}(\models_{2}\cap~(\mathbf{M}\times\{\{\alpha\}\}))$. Now, $m\in\pi_{1}(\models_{1}\cap~(\mathbf{M}\times\{\Sigma\}))$ implies that $m\models_{1}\Sigma$. Since $\mathfrak{S}$ is strongly granular (and hence granular at the 1st component), $m\models_{1}\Sigma$ implies that $m\models_{1}\Gamma$. Consequently, $m\in\pi_{1}(\models_{1}\cap~(\mathbf{M}\times\{\Gamma\}))$ but $m\notin\pi_{1}(\models_{2}\cap~(\mathbf{M}\times\{\{\alpha\}\}))$, i.e., $\pi_{1}(\models_{1}\cap~(\mathbf{M}\times\{\Gamma\}))\,\not\subseteq\,\pi_{1}(\models_{2}\cap~(\mathbf{M}\times\{\{\alpha\}\}))$. ∎
    By Theorem 3.8, every $q$-consequence operator has an adequate functionally $3$-valued semantics. However, not every $q$-consequence operator is inferentially $3$-valued. For example, every logical structure induced by a reflexive $q$-consequence operator is of Tarski-type, and it is inferentially 2-valued. However, as shown below, every non-reflexive $q$-consequence operator is inferentially 3-valued.
    Theorem 3.10 (Inferential Irreducibility Theorem for $q$-type).
    Every logical structure induced by a non-reflexive $q$-consequence operator is inferentially 3-valued.
    Proof.
    Let $(\mathscr{L},W)$ be a logical structure induced by a non-reflexive $q$-consequence operator. Note that, by Theorem 3.8, it is functionally $3$-valued. Suppose that it is not inferentially $3$-valued. So, there exists some functionally $2$-valued semantics for $\mathscr{L}$, say $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S_{q},\subseteq,\mathcal{P}(\mathscr{L}))$ such that $W=W_{\mathfrak{S}}$. Since $\mathfrak{S}$ is functionally $2$-valued, without loss of generality, we may assume that $\mathbf{M}\subseteq\{0,1\}^{\mathscr{L}}$.
    So, for each $i\in\{1,2\}$, there exists $\emptyset\subsetneq D_{i}\subsetneq\{0,1\}$ such that $D_{1}\cup D_{2}\subsetneq\{0,1\}$. Since $\emptyset\subsetneq D_{1}\cup D_{2}\subsetneq\{0,1\}$, either $D_{1}\cup D_{2}=\{1\}$ or $D_{1}\cup D_{2}=\{0\}$. Without loss of generality, let us assume that $D_{1}\cup D_{2}=\{1\}$. Consequently, since $\emptyset\subsetneq D_{i}$ for all $i\in\{1,2\}$, it follows that $D_{1}=\{1\}=D_{2}$.
    Thus, for all $m\in\mathbf{M},\Gamma\subseteq\mathscr{L}$, $m\models_{1}\Gamma$ implies that $m(\Gamma)\subseteq\{1\}$. Similarly, for all $m\in\mathbf{M},\Gamma\subseteq\mathscr{L}$, $m\models_{2}\Gamma$ implies that $m(\Gamma)\subseteq\{1\}$. Therefore, $\models_{1}\,=\,\models_{2}$.
    We now claim that $W_{\mathfrak{S}}$ is reflexive. To prove this, we first show that $\alpha\in W_{\mathfrak{S}}(\{\alpha\})$ for all $\alpha\in\mathscr{L}$. Suppose the contrary. Then there exists $\alpha\in\mathscr{L}$ such that $\alpha\notin W_{\mathfrak{S}}(\{\alpha\})$. But this would mean that there exists $m\in\mathbf{M}$ such that $m\models_{1}\{\alpha\}$ such that $m\not\models_{2}\{\alpha\}$ - a contradiction since $\models_{1}\,=\,\models_{2}$. Thus $\alpha\in W_{\mathfrak{S}}(\{\alpha\})$ for all $\alpha\in\mathscr{L}$. Now, let $\Gamma\subseteq\mathscr{L}$ and $\beta\in\Gamma$, i.e., $\{\beta\}\subseteq\Gamma$. Since $W=W_{\mathfrak{S}}$ and $W$ is monotonic, $W_{\mathfrak{S}}$ is monotonic as well. This implies that $W_{\mathfrak{S}}(\{\beta\})\subseteq W_{\mathfrak{S}}(\Gamma)$. Since $\beta\in W_{\mathfrak{S}}(\{\beta\})$, we must have $\beta\in W_{\mathfrak{S}}(\Gamma)$. So $\beta\in W_{\mathfrak{S}}(\Gamma)$ for all $\beta\in\Gamma$. Consequently, $W_{\mathfrak{S}}$, which implies that $W$ is reflexive as well - a contradiction. This establishes the theorem. ∎
    Corollary 3.11 (Inferential Characterisation of the Class of $q$-type).
    The class of $q$-type logical structures is inferentially 3-valued.
  • 3.2 $p$-consequence Operators
    $p$-consequence operators were introduced by Frankowski in order to formalise the notion of ‘plausible inferences’ in [18, p. 1] as follows.
    The key intuition behind plausibility in the sense of Ajdukiewicz [1] is that it admits reasonings wherein the degree of strength of the conclusion (i.e. the conviction it is true) is smaller then that of the premises. In consequence in the formal description of plausible inference, the standard notions of consequence and rule of inference cannot be countenanced as entirely suitable.
    Following is the formal definition of a $p$-consequence operator.
    Definition 3.12 ($p$-consequence Operator).
    A $p$-consequence operator on $\mathscr{L}$ is a function $W:\mathcal{P}(\mathscr{L})\to\mathcal{P}(\mathscr{L})$ satisfying the following properties.
    1. (i)
      For all $\Gamma\subseteq\mathscr{L}$, $\Gamma\subseteq W(\Gamma)$. (Reflexivity)
    2. (ii)
      For all $\Gamma\cup\Sigma\subseteq\mathscr{L}$, if $\Gamma\subseteq\Sigma$ then $W(\Gamma)\subseteq W(\Sigma)$. (Monotonicity)
    A logical structure $(\mathscr{L},W)$ is said to be of p-type if $W$ is a $p$-consequence operator on $\mathscr{L}$.
    The following theorem provides another criteria for characterising $p$-consequence operators, which we will need in a later theorem.
    Theorem 3.13 (Characterisation of $p$-type).
    Let $(\mathscr{L},W)$ be a logical structure. Then, $W$ is a $p$-consequence operator iff for all $\Gamma\cup\Sigma\subseteq\mathscr{L}$, $\Gamma\subseteq\Sigma$ implies that $\Gamma\cup W(\Gamma)\subseteq W(\Sigma)$.
    Proof.
    Suppose $W$ is a $p$-consequence operator. Let $\Gamma\cup\Sigma\subseteq\mathscr{L}$ such that $\Gamma\subseteq\Sigma$. Since $W$ is reflexive, $\Sigma\subseteq W(\Sigma)$, which implies that $\Gamma\subseteq W(\Sigma)$. Moreover as $W$ is monotonic, $\Gamma\subseteq\Sigma$ implies that $W(\Gamma)\subseteq W(\Sigma)$ as well. Consequently, $\Gamma\cup W(\Gamma)\subseteq W(\Sigma)$ for all $\Gamma\cup\Sigma\subseteq\mathscr{L}$.
    Conversely, suppose that for all $\Gamma\cup\Sigma\subseteq\mathscr{L}$, $\Gamma\subseteq\Sigma$ implies $\Gamma\cup W(\Gamma)\subseteq W(\Sigma)$. Let $\Gamma\subseteq\mathscr{L}$. Since $\Gamma\subseteq\Gamma$, $\Gamma\cup W(\Gamma)\subseteq W(\Gamma)$, which implies that $\Gamma\subseteq W(\Gamma)$. Thus, $W$ is reflexive.
    Let $\Gamma\cup\Sigma\subseteq\mathscr{L}$ such that $\Gamma\subseteq\Sigma$. Then, $\Gamma\cup W(\Gamma)\subseteq W(\Sigma)$ which implies that $W(\Gamma)\subseteq W(\Sigma)$. So $W$ is monotonic as well.
    Hence, $W$ is a $p$-consequence operator. ∎
    Theorem 3.14 (Representation Theorem for $p$-type (Part I)).
    Let $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S_{p},\mathcal{P}(\mathscr{L}))$ be a strongly granular semantics for $\mathscr{L}$ such that $S_{p}=\{(\models_{1},\models_{2})\}$ and $\models_{1}\,\subseteq\,\models_{2}$. Then, $W_{\mathfrak{S}}$ is a $p$-consequence operator.
    Proof.
    Let $\Gamma\cup\Sigma\subseteq\mathscr{L}$ such that $\Gamma\subseteq\Sigma$ but $\Gamma\cup W_{\mathfrak{S}}(\Gamma)\not\subseteq W_{\mathfrak{S}}(\Sigma)$. This implies that there exists $\alpha\in\Gamma\cup W_{\mathfrak{S}}(\Gamma)$ such that $\alpha\notin W_{\mathfrak{S}}(\Sigma)$.
    Since $\alpha\notin W_{\mathfrak{S}}(\Sigma)$, there exists $m\in\mathbf{M}$ such that $m\models_{1}\Sigma$ but $m\not\models_{2}\{\alpha\}$. Since $\mathfrak{S}$ is strongly granular (and hence granular at the 1st component), $\Gamma\subseteq\Sigma$ and $m\models_{1}\Sigma$, it follows that $m\models_{1}\Gamma$. Since $\alpha\in\Gamma\cup W_{\mathfrak{S}}(\Gamma)$, either $\alpha\in\Gamma$ or $\alpha\in W_{\mathfrak{S}}(\Gamma)$.
    If $\alpha\in\Gamma$, then as $m\models_{1}\Gamma$ and $\mathfrak{S}$ is strongly granular, $m\models_{1}\{\alpha\}$ as well. Since $\models_{1}\,\subseteq\,\models_{2}$, this implies $m\models_{2}\{\alpha\}$ - a contradiction. If, on the other hand, $\alpha\in W_{\mathfrak{S}}(\Gamma)$, then since $m\models_{1}\Gamma$, we have $m\models_{2}\{\alpha\}$ - a contradiction again. Thus $\Gamma\cup W_{\mathfrak{S}}(\Gamma)\subseteq W_{\mathfrak{S}}(\Sigma)$. Therefore, by the previous theorem, $W$ is a $p$-consequence operator. ∎
    It has been proved in [19, Theorem 5] that every $p$-consequence operator can be given a 3-valued description. In our language this is equivalent to stating that every $p$-type logical structure has an adequate functionally $3$-valued semantics. Below we give a proof of the same result mainly for completeness.
    Theorem 3.15 (Representation Theorem for $p$-type (Part II)).
    Let $(\mathscr{L},W)$ be a $p$-type logical structure and $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S_{p},\mathcal{P}(\mathscr{L}))$ a semantics for $\mathscr{L}$ satisfying the following properties.
    1. $\bullet$
      $\mathbf{M}=\{\mu_{\Sigma}\in\{0,1,2\}^{\mathscr{L}}:\Sigma\subseteq\mathscr{L}\}$ where
      $\mu_{\Sigma}(\beta)=\begin{cases}0&\text{if}~\beta\notin W(\Sigma)\\ 1&\text{if}~\beta\in\Sigma\\ 2&\text{if}~\beta\in W(\Sigma)\setminus\Sigma\end{cases}$
    2. $\bullet$
      For all $m\in\mathbf{M},\Gamma\subseteq\mathscr{L}$,
      $m\models_{1}\Gamma~\text{iff}~m(\Gamma)\subseteq\{1\}$
      $m\models_{2}\Gamma~\text{iff}~m(\Gamma)\subseteq\{1,2\}$
    3. $\bullet$
      $S_{p}$ is as defined in the previous theorem.
    Then, $W=W_{\mathfrak{S}}$.
    Proof.
    Let for some $\Gamma\subseteq\mathscr{L}$, $W(\Gamma)\not\subseteq W_{\mathfrak{S}}(\Gamma)$. Then, there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W_{\mathfrak{S}}(\Gamma)$. So, there exists $m\in\textbf{M}$ such that $m\models_{1}\Gamma$ but $m\not\models_{2}\{\alpha\}$. Since $m\in\textbf{M}$, there exists $\Sigma\subseteq\mathscr{L}$ such that $m=\mu_{\Sigma}$. Now, $\mu_{\Sigma}\models_{1}\Gamma$ implies that $\mu_{\Sigma}(\Gamma)\subseteq\{1\}$, i.e., $\Gamma\subseteq\Sigma$. Since $W$ is monotonic, $W(\Gamma)\subseteq W(\Sigma)$. So, $\alpha\in W(\Sigma)$, and hence $\mu_{\Sigma}\models_{2}\{\alpha\}$, i.e., $m\models_{2}\{\alpha\}$ - a contradiction. Consequently, $W(\Gamma)\subseteq W_{\mathfrak{S}}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$.
    Next, let $\Gamma\subseteq\mathscr{L}$ and $\alpha\in W_{\mathfrak{S}}(\Gamma)$. Now, $\mu_{\Gamma}(\Gamma)\subseteq\{1\}$ and so $\mu_{\Gamma}\models_{1}\Gamma$. As $\alpha\in W_{\mathfrak{S}}(\Gamma)$, $\mu_{\Gamma}\models_{2}\{\alpha\}$. Thus, $\mu_{\Gamma}(\{\alpha\})\subseteq\{1,2\}$. If $\mu_{\Gamma}(\alpha)=1$ then $\alpha\in\Gamma$, and since $W$ is reflexive, $\alpha\in W(\Gamma)$ as well. If, on the other hand, $\mu_{\Gamma}(\alpha)=2$ then also $\alpha\in W(\Gamma)$. Thus, in either case, $\alpha\in W(\Gamma)$. Hence, $W_{\mathfrak{S}}(\Gamma)\subseteq W(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$. This completes the proof. ∎
    Theorem 3.16 (Representation Theorem for $p$-type (Part III)).
    Let $(\mathscr{L},W)$ be a $p$-type logical structure and $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S,\mathcal{P}(\mathscr{L}))$ a strongly granular semantics for $\mathscr{L}$ such that $S=\{(\models_{1},\models_{2})\}$ and $W=W_{\mathfrak{S}}$. Then $\models_{1}\,\subseteq\,\models_{2}$.
    Proof.
    Let $\alpha\in\mathscr{L}$ and $m\in\mathbf{M}$ be such that $m\models_{1}\{\alpha\}$. If possible, let $m\not\models_{2}\{\alpha\}$. But then $\alpha\notin W_{\mathfrak{S}}(\{\alpha\})$ which implies that $\alpha\notin W(\{\alpha\})$ as $W=W_{\mathfrak{S}}$. This, however, is a contradiction since $W$ is reflexive. Thus, for all $\alpha\in\mathscr{L},m\in\mathbf{M}$, $m\models_{1}\{\alpha\}$ implies that $m\models_{2}\{\alpha\}$.
    Now, let $\Gamma\subseteq\mathscr{L}$ be such that for some $n\in\mathbf{M}$, $n\models_{1}\Gamma$. Then, as $\mathfrak{S}$ is strongly granular (and hence granular at the first component) $n\models_{1}\{\gamma\}$ for all $\gamma\in\Gamma$. So, by the above argument $n\models_{2}\{\gamma\}$ for all $\gamma\in\Gamma$. Then again, as $\mathfrak{S}$ is strongly granular (and hence granular at the second component) $n\models_{2}\Gamma$. Hence, $\models_{1}\,\subseteq\,\models_{2}$. ∎
    By Theorem 3.16 every $p$-consequence operator has an adequate functionally $3$-valued semantics. However, not every $p$-type logical structure is inferentially $3$-valued.
    Given a logical structure $(\mathscr{L},W)$, $W$ is said to be idempotent if $W=W\circ W$. We now show that every logical structure induced by a non-idempotent $p$-consequence operator is inferentially $3$-valued.
    Theorem 3.17 (Inferential Irreducibility Theorem for $p$-type).
    Every logical structure induced by a non-idempotent $p$-consequence operator is inferentially 3-valued.
    Proof.
    Let $(\mathscr{L},W)$ be a logical structure induced by a non-idempotent $p$-consequence operator. Suppose that it is not inferentially $3$-valued. Then, there exists some functionally $2$-valued semantics for $\mathscr{L}$, say $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S_{p},\mathcal{P}(\mathscr{L}))$ where $S_{p}=\{(\models_{1},\models_{2})\}$ and $\models_{1}\,\subseteq\,\models_{2}$, such that $W=W_{\mathfrak{S}}$.
    Using an argument similar to that in Theorem 3.10, we can show that for all $m\in\mathbf{M},\Gamma\subseteq\mathscr{L}$,
    $m\models_{1}\Gamma\iff m(\Gamma)\subseteq\{1\}\iff m\models_{2}\Gamma$
    In other words, $\models_{1}\,=\,\models_{2}$. Let $\models_{1}\,=\,\models_{2}\,=\,\models$.
    By Remark 2.15, every functionally 2-valued semantics is strongly granular. Now, let $\Sigma\subseteq\mathscr{L}$ such that $W_{\mathfrak{S}}(W(\Sigma))\not\subseteq W_{\mathfrak{S}}(\Sigma)$. Then, there exists $\alpha\in W_{\mathfrak{S}}(W(\Sigma))$ such that $\alpha\notin W_{\mathfrak{S}}(\Sigma)$. So, there exists $m\in\mathbf{M}$ such that $m\models\Sigma$ such that $m\not\models\{\alpha\}$. Observe that, for all $\beta\in W(\Sigma)$, $m\models\{\beta\}$. Consequently, $m\models W(\Sigma)$. Now, as $m\not\models\{\alpha\}$, $\alpha\notin W_{\mathfrak{S}}(W(\Sigma))$ - a contradiction. Hence $W_{\mathfrak{S}}(W(\Sigma))\subseteq W_{\mathfrak{S}}(\Sigma)$.
    Since $W=W_{\mathfrak{S}}$, this shows that $W(W(\Sigma))\subseteq W(\Sigma)$. Now, by reflexivity of $W$, we have $W(\Sigma)\subseteq W(W(\Sigma))$. This shows that $W(W(\Sigma))=W(\Sigma)$ for all $\Sigma\subseteq\mathscr{L}$. Thus, $W$ is idempotent - a contradiction. This completes the proof. ∎
    Corollary 3.18 (Inferential Characterisation of the Class of $p$-type).
    The class of $p$-type logical structures is inferentially 3-valued.
    3.3 $s_{\kappa}$-consequence Operators
    Let $(\mathscr{L},W)$ be a logical structure such that $W$ is reflexive. We now discuss a family of non-reflexive monotonic consequence operators. Given a logical structure $(\mathscr{L},W)$, a nonempty set $\Gamma\subseteq\mathscr{L}$, is said to satisfy anti-reflexivity if $\Gamma\subseteq\mathscr{L}\setminus W(\Gamma)$, i.e., $\Gamma\cap W(\Gamma)=\emptyset$. We will say that $(\mathscr{L},W)$ is anti-reflexive if for all $\Gamma\subseteq\mathscr{L}$, $\Gamma$ satisfies anti-reflexivity. We note that there exists only one monotonic logical structure $(\mathscr{L},W)$ such that $\Gamma$ is anti-reflexive for all $\Gamma\subseteq\mathscr{L}$. More precisely, we have the following theorem.
    Theorem 3.19.
    A monotonic logical structure $(\mathscr{L},W)$ with $\mathscr{L}\neq\emptyset$ is anti-reflexive iff for all nonempty $\Gamma\subseteq\mathscr{L}$, $W(\Gamma)=\emptyset$.
    Proof.
    Let $(\mathscr{L},W)$ be a monotonic logical structure such that for all nonempty $\Gamma\subseteq\mathscr{L}$, $W(\Gamma)=\emptyset$. Then, in particular, $W(\mathscr{L})=\emptyset$. Since $W$ is monotonic, for any $\Gamma\subseteq\mathscr{L}$, $W(\Gamma)=\emptyset\subseteq\emptyset=W(\mathscr{L})$. Therefore, for any nonempty $\Gamma\subseteq\mathscr{L}$, $\Gamma\subseteq\mathscr{L}\setminus W(\Gamma)=\mathscr{L}\setminus\emptyset=\mathscr{L}$. Consequently, $W$ is anti-reflexive as well.
    Conversely, suppose that $(\mathscr{L},W)$ is anti-reflexive. Now, as $\mathscr{L}\neq\emptyset$, by anti-reflexivity, $\mathscr{L}\cap W(\mathscr{L})=\emptyset$, which implies $W(\mathscr{L})=\emptyset$. Then, for any nonempty $\Gamma\subseteq\mathscr{L}$, $\emptyset\subseteq W(\Gamma)\subseteq W(\mathscr{L})=\emptyset$, i.e., $W(\Gamma)=\emptyset$ by monotonicity, and we are done. ∎
    The above theorem shows that the concept of ‘global’ anti-reflexivity in the presence of monotonicity is too strong. This, however, is not the case for non-monotonic logical structures (see, e.g., [2, 8]). This is also not the case if one considers certain ‘relativised’ versions of anti-reflexivity.
    Definition 3.20 (Internally $\kappa$ Set).
    Given a set $\mathscr{L}$, a cardinal $\kappa$, and $\mathsf{K}\subseteq\mathcal{P}(\mathscr{L})$, we say that $\mathsf{K}$ is internally $\kappa$ if for all $\Gamma\in\mathsf{K}$, $|\Gamma|\geq\kappa$ and there exists $\Sigma\in\mathsf{K}$ such that $|\Sigma|=\kappa$.
    Definition 3.21 ($\textsc{Anti-reflexivity}^{\mathsf{K}}_{\kappa}$).
    Given a logical structure $(\mathscr{L},W)$, a cardinal $\kappa$ and $\mathsf{K}\subseteq\mathcal{P}(\mathscr{L})$, $W$ is said to be $\textit{anti-reflexive}^{\mathsf{K}}_{\kappa}$ (anti-reflexive relative to $\mathsf{K}$ that is internally $\kappa$) or to satisfy $\textit{anti-reflexivity}^{\mathsf{K}}_{\kappa}$ if for every $\Gamma\in\mathsf{K}$, $\Gamma$ satisfies anti-reflexivity, i.e., $\Gamma\subseteq\mathscr{L}\setminus W(\Gamma)$ for all $\Gamma\in\mathsf{K}$. In this case, $(\mathscr{L},W)$ is also said to $\textit{be anti-reflexive}^{\mathsf{K}}_{\kappa}$.
    Definition 3.22 ($s^{\mathsf{K}}_{\kappa}$-consequence Operator).
    Let $\mathscr{L}$ be a set, $W:\mathcal{P}(\mathscr{L})\to\mathcal{P}(\mathscr{L})$ and $\kappa$ a cardinal. Then $W$ is said to be an $s^{\mathsf{K}}_{\kappa}$-consequence operator on $\mathscr{L}$ if it is monotonic and $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$. A logical structure $(\mathscr{L},W)$ is said to be of $s^{\mathsf{K}}_{\kappa}\textit{-type}$ if $W$ is an $s^{\mathsf{K}}_{\kappa}$-consequence on $\mathscr{L}$.
    $W$ is said to be an $s_{\kappa}$-consequence operator on $\mathscr{L}$ if $W$ is an $s^{\mathsf{K}}_{\kappa}$-consequence operator on $\mathscr{L}$ for some $\mathsf{K}\subseteq\mathscr{L}$ that is internally $\kappa$. A logical structure $(\mathscr{L},W)$ is said to be of $s_{\kappa}\textit{-type}$ if $W$ is an $s_{\kappa}$-consequence operator on $\mathscr{L}$.
    We now list a few properties of $s_{\kappa}$-consequence operators.
    Theorem 3.23.
    Let $(\mathscr{L},W)$ be an $s^{\mathsf{K}}_{\kappa}$-type logical structure where $\mathscr{L}\neq\emptyset$. Then, for all $\Gamma\in\mathsf{K}$, the following statements hold.
    1. (1)
      $W(\Gamma)\subseteq W(\mathscr{L}\setminus W(\Gamma))$. The reverse inclusion holds if $\mathscr{L}\setminus W(\Gamma)\in\mathsf{K}$ as well.
    2. (2)
      If $\mathscr{L}\setminus\Gamma\in\mathsf{K}$, then $W(\Gamma)\cap W(\mathscr{L}\setminus\Gamma)=\emptyset$.
    3. (3)
      If $\Sigma\in\mathsf{K}$ such that $\Gamma\cap\Sigma\neq\emptyset$, then $W(\Sigma)\neq\Gamma$
    Proof.
    (1) Let $\Gamma\in\mathsf{K}$. Then $\Gamma\subseteq\mathscr{L}\setminus W(\Gamma)$ by $\text{anti-reflexivity}^{\mathsf{K}}_{\kappa}$ of $W$. So, by monotonicity, $W(\Gamma)\subseteq W(\mathscr{L}\setminus W(\Gamma))$.
    Now, suppose that $\mathscr{L}\setminus W(\Gamma)\in\mathsf{K}$ as well. Then, for any $\alpha\notin W(\Gamma)$, $\alpha\in\mathscr{L}\setminus W(\Gamma)$, which implies that $\alpha\notin W(\mathscr{L}\setminus W(\Gamma))$ as $W$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$. Thus, $W(\mathscr{L}\setminus W(\Gamma))\subseteq W(\Gamma)$.
    (2) Let $\Gamma\in\mathsf{K}$ such that $\mathscr{L}\setminus\Gamma\in\mathsf{K}$. Then, for any $\alpha\in W(\Gamma)$,
    $\alpha\notin\mathscr{L}\setminus W(\Gamma)$ $\implies\alpha\notin\Gamma$ $(\text{since}~W~\text{is}~\text{anti-reflexive}^{\mathsf{K}}_{\kappa})$
    $\implies\alpha\in\mathscr{L}\setminus\Gamma$
    $\implies\alpha\in\mathscr{L}\setminus W(\mathscr{L}\setminus\Gamma)$ $(\text{since}~\mathscr{L}\setminus\Gamma\in\mathsf{K}~\text{and}~W~\text{is}~\text{anti-reflexive}^{\mathsf{K}}_{\kappa})$
    $\implies\alpha\notin W(\mathscr{L}\setminus\Gamma)$
    Thus, $W(\Gamma)\cap W(\mathscr{L}\setminus\Gamma)=\emptyset$.
    (3) Suppose there exists $\Gamma\subseteq\mathscr{L}$ and $\Sigma\in\mathsf{K}$ with $\Gamma\cap\Sigma\neq\emptyset$ such that $W(\Sigma)=\Gamma$. Now, since $W$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$, we have $\Gamma=W(\Sigma)\subseteq\mathscr{L}\setminus\Sigma$, which implies that $\Gamma\cap\Sigma=\emptyset$ - a contradiction. Hence $W(\Sigma)\neq\Gamma$. ∎
    Example 3.24.
    Let, $(\mathscr{L},W)$ be a logical structure, where $\mathscr{L}=\mathbb{N}\times\mathbb{N}$ and $W:\mathcal{P}(\mathscr{L})\to\mathcal{P}(\mathscr{L})$ is defined as follows.
    $W(\Gamma)=\begin{cases}\{(n,m):(m,n)\in\Gamma\}&\text{if}~\Gamma\neq\emptyset\\ \emptyset&\text{otherwise}\end{cases}$
    for any $\Gamma\subseteq\mathscr{L}$.
    Let $\Phi=\{(2n+1,2n+2):n\in\mathbb{N}\}$ and $\mathsf{K}=\mathcal{P}(\Phi)\setminus\{\emptyset\}$. It is easy to see that $\mathsf{K}$ is internally 1. We now claim that for all $\Gamma\in\mathsf{K}$, $\Gamma\subseteq\mathscr{L}\setminus W(\Gamma)$.
    Suppose the contrary, i.e., there exists $\Gamma\in\mathsf{K}$, $\Gamma\cap\mathscr{L}\setminus W(\Gamma)\neq\emptyset$. Let $(a,b)\in\Gamma\cap W(\Gamma)$. Now, $(a,b)\in W(\Gamma)$ implies that $(b,a)\in\Gamma\subseteq\Phi$ - an impossibility! Thus, $\Gamma\cap\mathscr{L}\setminus W(\Gamma)=\emptyset$ for all $\Gamma\in\mathsf{K}$.
    Hence $(\mathscr{L},W)$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$.
    Next, let $\Gamma\cup\Sigma\subseteq\mathscr{L}$ such that $\Gamma\subseteq\Sigma$. If $\Gamma=\emptyset$, then $W(\Gamma)=\emptyset$, and hence $W(\Gamma)\subseteq W(\Sigma)$.
    Suppose that $\Gamma\neq\emptyset$. So $W(\Gamma)\neq\emptyset$. Then,
    $(n,m)\in W(\Gamma)$ $\implies(m,n)\in\Gamma$ $\text{(by the definition of}~W)$
    $\implies(m,n)\in\Sigma$ $\text{(since}~\Gamma\subseteq\Sigma)$
    $\implies(n,m)\in W(\Sigma)$ $\text{(by the definition of}~W)$
    So $W$ is monotonic as well. Thus, $W$ is an $s^{\mathsf{K}}_{1}$-consequence operator.
    We now show that given any infinite set $\mathscr{L}$ and $0<\lambda<|\mathscr{L}|$, there exists a logical structure $(\mathscr{L},W)$ and $\emptyset\subsetneq\mathsf{K}\subseteq\mathcal{P}(\mathscr{L})$ such that $W$ is an $s^{\mathsf{K}}_{\kappa}$-consequence operator. But for this, we need the following result about cardinal numbers.
    Theorem 3.25.
    Given a set $X$, and a cardinal $\kappa$ such that $|X|=\kappa$, there exists $Y\subseteq X$ such that $|Y|=\lambda$ for all $\lambda\leq\kappa$.
    Example 3.26.
    Let $\kappa$ be an infinite cardinal and $\mathscr{L}$ a set such that $|\mathscr{L}|=\kappa$.
    Then, $|\mathscr{L}\times\{0\}|=|\mathscr{L}\times\{1\}|=\kappa$. So, there exist a bijection $\sigma:\mathscr{L}\to(\mathscr{L}\times\{0\})\cup(\mathscr{L}\times\{1\})$. Let $X=\sigma^{-1}(\mathscr{L}\times\{0\})$ and $Y=\sigma^{-1}(\mathscr{L}\times\{1\})$. Now,
    $X\cap Y=\sigma^{-1}(\mathscr{L}\times\{0\})\cap\sigma^{-1}(\mathscr{L}\times\{1\})=\sigma^{-1}((\mathscr{L}\times\{0\})\cap(\mathscr{L}\times\{1\}))=\sigma^{-1}(\emptyset)=\emptyset$
    and
    $X\cup Y=\sigma^{-1}(\mathscr{L}\times\{0\})\cup\sigma^{-1}(\mathscr{L}\times\{1\})=\sigma^{-1}((\mathscr{L}\times\{0\})\cup(\mathscr{L}\times\{1\}))=\mathscr{L}.$
    Moreover, $|X|=|Y|=\kappa$.
    We now define $f:\mathscr{L}\to\mathscr{L}$ as follows.
    $f(z)=\begin{cases}\sigma^{-1}(a,1)&\text{if}~\sigma(z)=(a,0)\\ \sigma^{-1}(a,0)&\text{if}~\sigma(z)=(a,1)\end{cases}$
    Well-definedness of $f$ follows from the fact that $\sigma$ is a bijection.
    Let us now choose any $0<\lambda<\kappa$ and define, $W_{\lambda}:\mathcal{P}(\mathscr{L})\to\mathcal{P}(\mathscr{L})$ as follows.
    $W_{\lambda}(\Gamma)=\begin{cases}f(\Gamma)&\text{if}~|\Gamma|\geq\lambda\\ \emptyset&\text{otherwise}\end{cases}$
    Finally, let $\mathsf{K}=\{\Gamma\subseteq\lambda:|\Gamma|\geq\lambda~\text{and}~\Gamma\subseteq X~\text{or}~\Gamma\subseteq Y\}$ and consider the logical structure $(\mathscr{L},W_{\lambda})$.
    We note that since $\lambda<\kappa$, by Theorem 3.25, there exists $\Delta\subseteq\mathscr{L}$ such that $|\Delta|=\lambda$. Moreover, for every $\Gamma\in\mathsf{K}$, $|\Gamma|\geq\lambda$. Therefore, $\mathsf{K}$ is internally $\lambda$.
    We now claim that for all $\Gamma\in\mathsf{K}$, $\Gamma\subseteq\mathscr{L}\setminus W_{\lambda}(\Gamma)$. Let $\Gamma\in\mathsf{K}$. Then, $|\Gamma|\geq\lambda$ and $\Gamma\subseteq X$ or $\Gamma\subseteq Y$. Without loss of generality, suppose that $\Gamma\subseteq X$. Then, $W_{\lambda}(\Gamma)=f(\Gamma)\subseteq Y$. Now, as $X\cap Y=\emptyset$, $\Gamma\cap W_{\lambda}(\Gamma)=\emptyset$, i.e., $\Gamma\subseteq\mathscr{L}\setminus W_{\lambda}(\Gamma)$. Thus $W_{\kappa}$ is $\text{anti-reflexive}^{\mathsf{K}}_{\lambda}$.
    Next, let $\Gamma\cup\Sigma\subseteq\mathscr{L}$ such that $\Gamma\subseteq\Sigma$. If $|\Gamma|<\lambda$ then $W_{\lambda}(\Gamma)=\emptyset$ which implies $W_{\lambda}(\Gamma)\subseteq W_{\lambda}(\Sigma)$. If, on the other hand, $|\Gamma|\geq\lambda$, then as $\Gamma\subseteq\Sigma$, $|\Sigma|\geq\lambda$ as well. So,
    $W_{\lambda}(\Gamma)=f(\Gamma)\subseteq f(\Sigma)=W_{\kappa}(\Sigma)$
    Therefore $W_{\lambda}$ is monotonic. Consequently, $W_{\lambda}$ is an $s^{\mathsf{K}}_{\lambda}$-consequence operator.
    Remark 3.27.
    Given a logical structure $(\mathscr{L},W)$, where $W$ is an $s^{\mathsf{K}}_{\kappa}$-consequence operator, it can never be a $p$-consequence operator unless $\mathsf{K}=\{\emptyset\}$ and $\kappa=0$. Consequently, it follows that, the class of $p$-type logical structures is disjoint from the class of $s_{\kappa}$-type logical structures for all cardinal $\kappa>0$.
    However, every logical structure $(\mathscr{L},W)$, induced by a non-reflexive monotonic consequence operator is of $s_{1}$-type, as shown below.
    Theorem 3.28.
    Every non-reflexive and monotonic logical structure is of $s_{1}$-type.
    Proof.
    Let $(\mathscr{L},W)$ be a non-reflexive and monotonic logical structure. Then, by non-reflexivity, there exists $\Gamma\subseteq\mathscr{L}$ such that $\Gamma\not\subseteq W(\Gamma)$. Let $\alpha\in\Gamma$ such that $\alpha\notin W(\Gamma)$. Consequently, since $W$ is monotonic, $\alpha\notin W(\{\alpha\})$ as well. It is now easy to see that $W$ is $\text{anti-reflexive}_{1}^{\{\alpha\}}$. Since $W$ is monotonic as well, $W$ is an $s^{\{\alpha\}}_{1}-$type logical structure, and hence $(\mathscr{L},W)$ is of $s_{1}$-type. ∎
    Remark 3.29.
    For each cardinal $\kappa$, the class of $s_{\kappa}$-type logical structures is incomparable to that of the class of $q$-type logical structures. In other words, some $s_{\kappa}$-type logical structures are not of $q$-type, and vice versa. For example, the $s_{\kappa}$-consequence operator discussed in Example 3.26 is not a $q$-consequence operator. To see this, note that
    $W_{\kappa}(X\cup W_{\kappa}(X))=W_{\kappa}(X\cup Y)=W_{\kappa}(\mathscr{L})=f(\mathscr{L})=\mathscr{L}\neq Y=f(X)=W_{\kappa}(X)$
    However, the $s_{\kappa}$-consequence operator discussed in Theorem 3.19 is a $q$-consequence operator.
    Theorem 3.30 (Representation Theorem for $s^{\mathsf{K}}_{\kappa}$-type (Part I)).
    Let $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S_{s^{\mathsf{K}}_{\kappa}},\mathcal{P}(\mathscr{L}))$ be a strongly granular semantics for $\mathscr{L}$ and $\mathsf{K}\subseteq\mathcal{P}(\mathscr{L})$ such that the following properties hold.
    1. (1)
      $\mathsf{K}$ is internally $\kappa$.
    2. (2)
      $S_{s^{\mathsf{K}}_{\kappa}}=\{(\models_{1},\models_{2})\}$ and for all $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$ with $\Gamma\in\mathsf{K}$ and $\alpha\in\Gamma$ there exists $m\in\mathbf{M}$ such that $m\models_{1}\Gamma$ but $m\not\models_{2}\{\alpha\}$; in other words,
      $\pi_{1}(\models_{1}\cap\,(\mathbf{M}\times\{\Gamma\}))\not\subseteq\,\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\alpha\}\}))$
      for all $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$ with $\Gamma\in\mathsf{K}$ and $\alpha\in\Gamma$. Here $\pi_{1}:\mathbf{M}\times\mathcal{P}(\mathscr{L})\to\mathbf{M}$ is defined as follows: $\pi_{1}(m,\Delta)=m$ for all $(m,\Delta)\in\mathbf{M}\times\mathcal{P}(\mathscr{L})$.
    Then, $W_{\mathfrak{S}}$ is an $s^{\mathsf{K}}_{\kappa}$-consequence operator.
    Proof.
    First, we show that $W_{\mathfrak{S}}$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$. Suppose the contrary, and let $\Gamma\in\mathsf{K}$ such that $\Gamma\cap W_{\mathfrak{S}}(\Gamma)\neq\emptyset$. Let $\alpha\in\Gamma$ such that $\alpha\in W_{\mathfrak{S}}(\Gamma)$. Now, since $\alpha\in\Gamma$ and $\Gamma\in\mathsf{K}$, by (2), there exists $m\in\mathbf{M}$ such that $m\models_{1}\Gamma$ but $m\not\models_{2}\{\alpha\}$. This, however, implies that $\alpha\notin W_{\mathfrak{S}}(\Gamma)$ - a contradiction. Hence, $W_{\mathfrak{S}}$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$. Since $\mathfrak{S}$ is strongly granular, $W_{\mathfrak{S}}$ is monotonic, by Theorem 3.1. Thus, $W_{\mathfrak{S}}$ is an $s^{\mathsf{K}}_{\kappa}$-consequence operator. ∎
    We now show that every monotonic logical structure induced by an $s^{\mathsf{K}}_{\kappa}$-consequence operator is functionally 3-valued.
    Theorem 3.31 (Representation Theorem for $s^{\mathsf{K}}_{\kappa}$-type (Part II)).
    Let $(\mathscr{L},W)$ be an $s^{\mathsf{K}}_{\kappa}$-type logical structure for some $\mathsf{K}\subseteq\mathscr{L}$ that is internally $\kappa$. Let $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S_{s^{\mathsf{K}}_{\kappa}},\mathcal{P}(\mathscr{L}))$ be a semantics for $\mathscr{L}$ satisfying the following properties.
    1. $\bullet$
      $\mathbf{M}=\{\mu_{\Sigma}\in\{0,1,2\}^{\mathscr{L}}:\Sigma\subseteq\mathscr{L}\}$ where
      $\mu_{\Sigma}(\beta)=\begin{cases}0&\text{if}~\beta\notin W(\Sigma)\cup\Sigma\\ 1&\text{if}~\beta\in\Sigma\\ 2&\text{if}~\beta\in W(\Sigma)\end{cases}$
    2. $\bullet$
      For all $m\in\mathbf{M},\Gamma\subseteq\mathscr{L}$,
      $m\models_{1}\Gamma~\text{iff}~m(\Gamma)\subseteq\{1\}$
      $m\models_{2}\Gamma~\text{iff}~m(\Gamma)\subseteq\{2\}$
    3. $\bullet$
      $S_{s^{\mathsf{K}}_{\kappa}}=\{(\models_{1},\models_{2})\}$ and for all $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$ with $\Gamma\in\mathsf{K}$ and $\alpha\in\Gamma$,
      $\pi_{1}(\models_{1}\cap\,(\mathbf{M}\times\{\Gamma\}))\not\subseteq\,\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\alpha\}\}))$
      where $\pi_{1}:\mathbf{M}\times\mathcal{P}(\mathscr{L})\to\mathbf{M}$ is as defined in the previous theorem.
    Then, $W=W_{\mathfrak{S}}$.
    Proof.
    Suppose, there exists $\Gamma\subseteq\mathscr{L}$ such that $W(\Gamma)\not\subseteq W_{\mathfrak{S}}(\Gamma)$. Then, there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W_{\mathfrak{S}}(\Gamma)$. Now, $\alpha\notin W_{\mathfrak{S}}(\Gamma)$ implies that there exists $m\in\mathbf{M}$ such that $m\models_{1}\Gamma$ but $m\not\models_{2}\{\alpha\}$. Since $m\in\mathbf{M}$, $m=\mu_{\Sigma}$ for some $\Sigma\subseteq\mathscr{L}$.
    Now,
    $m\models_{1}\Gamma$ $\implies\mu_{\Sigma}\models_{1}\Gamma$
    $\implies\mu_{\Sigma}(\Gamma)\subseteq\{1\}$
    $\implies\Gamma\subseteq\Sigma$
    $\implies W(\Gamma)\subseteq W(\Sigma)$ $\text{(since}~W~\text{is monotonic)}$
    $\implies\alpha\in W(\Sigma)$ $\text{(since}~\alpha\in W(\Gamma))$
    On the other hand,
    $m\not\models_{2}\{\alpha\}$ $\implies\mu_{\Sigma}\not\models_{2}\{\alpha\}$
    $\implies\mu_{\Sigma}(\alpha)\neq 2$
    $\implies\alpha\notin W(\Sigma)$
    This is a contradiction. Thus $W(\Gamma)\subseteq W_{\mathfrak{S}}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$.
    Next, suppose that $W_{\mathfrak{S}}(\Gamma)\not\subseteq W(\Gamma)$. Then, there exists $\alpha\in W_{\mathfrak{S}}(\Gamma)$ such that $\alpha\notin W(\Gamma)$. Now, as $\Gamma\subseteq\Gamma$, $\mu_{\Gamma}(\Gamma)\subseteq\{1\}$, i.e., $\mu_{\Gamma}\models_{1}\Gamma$. However, since $\alpha\notin W(\Gamma)$, $\mu_{\Gamma}(\alpha)\neq 2$, which implies that $\mu_{\Gamma}\not\models_{2}\{\alpha\}$ and hence $\alpha\notin W_{\mathfrak{S}}(\Gamma)$ - a contradiction. Thus $W_{\mathfrak{S}}(\Gamma)\subseteq W(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$. Hence $W=W_{\mathfrak{S}}$. ∎
    Theorem 3.32 (Representation Theorem for $s^{\mathsf{K}}_{\kappa}$-type (Part III)).
    Let $(\mathscr{L},W)$ be an $s^{\mathsf{K}}_{\kappa}$-type logical for some $\mathsf{K}\subseteq\mathscr{L}$ that is internally $\kappa$. Let $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S,\mathcal{P}(\mathscr{L}))$ be a strongly granular semantics for $\mathscr{L}$ where $S=\{(\models_{1},\models_{2})\}$ and $W=W_{\mathfrak{S}}$. Then, for all $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$ with $\Gamma\in\mathsf{K}$ and $\alpha\in\Gamma$, we have,
    $\pi_{1}(\models_{1}\cap\,(\mathbf{M}\times\{\Gamma\}))\not\subseteq\,\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\alpha\}\}))$
    where $\pi_{1}:\mathbf{M}\times\mathcal{P}(\mathscr{L})\to\mathbf{M}$ is as defined in the above theorem.
    Proof.
    Suppose the contrary and let $\Gamma\in\mathsf{K}$ and $\alpha\in\Gamma$ such that $\pi_{1}(\models_{1}\cap\,(\mathbf{M}\times\{\Gamma\}))\subseteq\,\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\alpha\}\}))$. Now, since $W$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$, $\Gamma\in\mathsf{K}$ and $\alpha\in\Gamma$, $\alpha\notin W(\Gamma)$. Then, as $W=W_{\mathfrak{S}}$, $\alpha\notin W_{\mathfrak{S}}(\Gamma)$ as well. So, there exists $m\in\mathbf{M}$ such that $m\models_{1}\Gamma$ but $m\not\models_{2}\{\alpha\}$. Now, $m\not\models_{2}\{\alpha\}$ implies that $m\notin\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\alpha\}\}))$, and since $\pi_{1}(\models_{1}\cap\,(\mathbf{M}\times\{\Gamma\}))\subseteq\,\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\alpha\}\}))$, it follows that $m\not\models_{1}\Gamma$ - a contradiction. Consequently, $\pi_{1}(\models_{1}\cap\,(\mathbf{M}\times\{\Gamma\}))\not\subseteq\,\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\alpha\}\}))$ for all $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$ with $\Gamma\in\mathsf{K}$ and $\alpha\in\Gamma$. ∎
    The following theorem establishes the inferential $3$-valuedness of $s^{\mathsf{K}}_{\kappa}$-consequence operators. The proof is similar to Theorem 3.10 and hence we omit it.
    Theorem 3.33 (Inferential Irreducibility Theorem for $s^{\mathsf{K}}_{\kappa}$-type).
    Every $s_{\kappa}$-type logical structure is inferentially 3-valued.
    Corollary 3.34 (Inferential Characterisation of the Class of $s_{\kappa}$-type).
    For each cardinal $\kappa$, the class of $s_{\kappa}$-type logical structures is inferentially 3-valued.
    3.4 $r_{\kappa}$-consequence Operators
    We know from Theorem 3.2 that monotonic logical structures, in general, are functionally at most $4$-valued. However, the monotonic logical structures discussed so far in the previous subsections are all functionally at most $3$-valued. In this section, we identify the class of inferentially 4-valued monotonic logical structures.
    Suppose $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S,\mathcal{P}(\mathscr{L}))$ where $S=\{(\models_{1},\models_{2})\}$ is a strongly granular semantics for $\mathscr{L}$ such that $(\mathscr{L},W_{\mathfrak{S}})$ is an inferentially $4$-valued monotonic logical structure. We know from Theorem 3.8, Theorem 3.15 and Theorem 3.31 that every $q$-, $p$- and $s_{\kappa}$-type logical structure is functionally $3$-valued, and hence cannot be inferentially $4$-valued. Therefore, $(\mathscr{L},W_{\mathfrak{S}})$ cannot be a $q$-, $p$- or $s_{\kappa}$-type logical structure. This implies, in particular, that $W_{\mathfrak{S}}$ cannot be a $q$-, $p$- or $s_{\kappa}$-consequence operator.
    Now, by Theorem 3.7, if $\models_{2}\,\subseteq\,\models_{1}$, then $W_{\mathfrak{S}}$ is a $q$-consequence operator. Similarly, if $\models_{2}\,\subseteq\,\models_{1}$, then by Theorem 3.14 $W_{\mathfrak{S}}$ is a $q$-consequence operator. Therefore, since $W_{\mathfrak{S}}$ is not a $q$- or $p$-consequence operator, $\models_{1}$ and $\models_{2}$ must be incomparable.
    Moreover, if there exists $\mathsf{K}\subseteq\mathcal{P}(\mathscr{L})$ that is internally $\kappa$ for some cardinal $\kappa$ such that for all $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$ with $\Gamma\in\mathsf{K}$ and $\alpha\in\Gamma$, $\pi_{1}(\models_{1}\cap\,(\mathbf{M}\times\{\Gamma\}))\not\subseteq\,\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\alpha\}\}))$, then by Theorem 3.30, $W_{\mathfrak{S}}$ would be an $s_{\kappa}$-consequence operator. Therefore, since $W_{\mathfrak{S}}$ is not an $s_{\kappa}$-consequence operator, for all $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$ with $|\Gamma|\geq\kappa$ there exists $\alpha\in\Gamma$ such that $\pi_{1}(\models_{1}\cap\,(\mathbf{M}\times\{\Gamma\}))\subseteq\,\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\alpha\}\}))$.
    Now, let $(\mathscr{L},W)$ be a monotonic logical structure which is inferentially $4$-valued. Then, by Theorem 3.2 there exists a strongly granular semantics $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S,\mathcal{P}(\mathscr{L}))$ for $\mathscr{L}$ such that $W=W_{\mathfrak{S}}$. So, by the above discussion, $W$ is not a $q$-, $p$- or $s_{\kappa}$-consequence operator.
    1. (a)
      Since $W$ is not a $q$-consequence operator, by Theorem 3.6(5) there exists $\Gamma\subseteq\mathscr{L}$ such that $W(\Gamma)\not\subseteq\Gamma$. Otherwise, for any $\Gamma\subseteq\mathscr{L}$, $W(\Gamma)\cup\Gamma\subseteq\Gamma$, and hence $W(\Gamma)\cup\Gamma=\Gamma$ which implies that $W(W(\Gamma)\cup\Gamma)=W(\Gamma)$. This, however, would imply that $W$ satisfies quasi-closure, and so is a $q$-consequence operator - a contradiction.
    2. (b)
      Since $W$ is monotonic, but not a $p$-consequence operator, it is not reflexive. So, there exists $\Gamma\subseteq\mathscr{L}$ such that $\Gamma\not\subseteq W(\Gamma)$.
    3. (c)
      Since $W$ is monotonic, but not an $s_{\kappa}$-consequence operator, it is not $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$ for any $\mathsf{K}\subseteq\mathcal{P}(\mathscr{L})$ that is internally $\kappa$. This is the same as saying that for all $\Gamma\subseteq\mathscr{L}$ with $|\Gamma|\geq\kappa$, $\Gamma\not\subseteq\mathscr{L}\setminus W(\Gamma)$.
    The above discussion motivates the following definition of $r_{\kappa}$-consequence operators.
    Definition 3.35 ($r_{\kappa}$-consequence Operator).
    Given a cardinal $\kappa$, an $r_{\kappa}$-consequence operator on $\mathscr{L}$ is a function $W:\mathcal{P}(\mathscr{L})\to\mathcal{P}(\mathscr{L})$ satisfying the following properties.
    1. (i)
      There exists $\Gamma\subseteq\mathscr{L}$ such that $\Gamma\not\subseteq W(\Gamma)$. (Non-reflexivity)
    2. (ii)
      There exists $\Gamma\subseteq\mathscr{L}$ such that $W(\Gamma)\not\subseteq\Gamma$. (Non-quasi-closure)
    3. (iii)
      For all $\Gamma\subseteq\mathscr{L}$, with $|\Gamma|\geq\kappa$, $\Gamma\not\subseteq\mathscr{L}\setminus W(\Gamma)$. ($\text{Non-anti-reflexivity}_{\kappa}$)
    4. (iv)
      For all $\Gamma\cup\Sigma\subseteq\mathscr{L}$, if $\Gamma\subseteq\Sigma$ then $W(\Gamma)\subseteq W(\Sigma)$. (Monotonicity)
    A logical structure $(\mathscr{L},W)$ is said to be of $r_{\kappa}\textit{-type}$ if $W$ is an $r_{\kappa}$-consequence operator on $\mathscr{L}$.
    Theorem 3.36.
    There exists no logical structure $(\mathscr{L},W)$ such that $W$ is an $r_{1}$-consequence operator.
    Proof.
    Suppose the contrary. Let $(\mathscr{L},W)$ be a logical structure such that $W$ is an $r_{1}$-consequence operator. Since $W$ is non-reflexive, there exists $\Gamma\subseteq\mathscr{L}$ such that $\Gamma\not\subseteq W(\Gamma)$. Therefore, there exists $\alpha\in\Gamma$ such that $\alpha\notin W(\Gamma)$. Since $W$ is monotonic, this implies that $\alpha\notin W(\{\alpha\})$ as well. However, $\alpha\notin W(\{\alpha\})$ implies that $\alpha\in\mathscr{L}\setminus W(\{\alpha\})$, i.e., $\{\alpha\}\subseteq\mathscr{L}\setminus W(\{\alpha\})$, contradicting the fact that $W$ satisfies $\text{non-anti-reflexivity}_{1}$. Thus $(\mathscr{L},W)$ cannot be of $r_{1}$-type. Since $(\mathscr{L},W)$ was chosen arbitrarily, this completes the proof. ∎
    However, there exists an $r_{\kappa}$-consequence operator for any $\kappa\geq 2$, as shown in the following example.
    Example 3.37.
    Suppose $\mathscr{L}$ is an infinite set, and $\kappa\geq 2$ be a cardinal such that $|\mathscr{L}|>\kappa$. Let $\alpha\in\mathscr{L}$ and $W:\mathcal{P}(\mathscr{L})\to\mathcal{P}(\mathscr{L})$ be defined as follows.
    $W(\Gamma)=\begin{cases}\{\alpha\}&\text{if}~|\Gamma|<\kappa\\ \Gamma\cup\{\alpha\}&\text{otherwise}\end{cases}$
    We note that, since $\mathscr{L}$ is infinite, there exists $\beta\in\mathscr{L}$ such that $\alpha\neq\beta$. Now, since $|\{\beta\}|=1<\kappa$, $W(\{\beta\})=\{\alpha\}$. Then, as $\{\beta\}\not\subseteq W(\{\beta\})=\{\alpha\}$, $W$ is non-reflexive. Since $|\emptyset|=0<\kappa$, $W(\emptyset)=\{\alpha\}\not\subseteq\emptyset$. So, $W$ satisfies non-quasi-closure as well.
    Next, let $\Gamma\subseteq\mathscr{L}$ such that $|\Gamma|\geq\kappa$. Then, $W(\Gamma)=\Gamma\cup\{\alpha\}$, and hence $\Gamma\not\subseteq\mathscr{L}\setminus W(\Gamma)$. Thus $W$ satisfies $\text{non-anti-reflexivity}_{\kappa}$.
    Finally, let $\Gamma\cup\Sigma\subseteq\mathscr{L}$ such that $\Gamma\subseteq\Sigma$. If $|\Gamma|\geq\kappa$, then as $\Sigma\supseteq\Gamma$, $|\Sigma|\geq\kappa$ as well. So, $W(\Gamma)=\Gamma\cup\{\alpha\}\subseteq\Sigma\cup\{\alpha\}=W(\Sigma)$. If, on the other hand, $|\Gamma|<\kappa$, then $W(\Gamma)=\{\alpha\}$. Now, irrespective of $\Sigma$, $\{\alpha\}\subseteq W(\Sigma)$. Thus, in all cases $W(\Gamma)\subseteq W(\Sigma)$, and so $W$ is an $r_{\kappa}$-consequence operator.
    Remark 3.38.
    It is worth pointing out that the $r_{\kappa}$-consequence operator $W$ in the above example is not an $r_{\lambda}$-consequence operator for any cardinal $\lambda$ such that $2\leq\lambda<\kappa$. To see this, let $\lambda$ be a cardinal such that $2\leq\lambda<\kappa$ and $\Gamma_{\lambda}$ a set of cardinality $\lambda$ such that $\alpha\notin\Gamma_{\lambda}$. The existence of such a $\Gamma_{\lambda}$ can be justified as follows. If $\lambda$ is infinite, then we choose any set $\Gamma$ of cardinality $\lambda$ (which exists by Theorem 3.25), and take $\Gamma\setminus\{\alpha\}$ to be $\Gamma_{\lambda}$. On the other hand, if $\lambda$ is finite, we consider any set $\Gamma$ of cardinality $\lambda$, $\beta\notin\Gamma$ (such a $\beta$ exists since $\mathscr{L}$ is infinite but $\Gamma$ is finite) and define $\Gamma_{\lambda}=(\Gamma\setminus\{\alpha\})\cup\{\beta\}$. Then, as $|\Gamma_{\lambda}|=\lambda<\kappa$, $W(\Gamma_{\lambda})=\{\alpha\}$, which implies that $\Gamma_{\lambda}\subseteq\mathscr{L}\setminus W(\Gamma_{\lambda})=\mathscr{L}\setminus\{\alpha\}$. This shows that $W$ does not satisfy $\text{non-anti-reflexivity}_{\lambda}$. Thus, $W$ is not an $r_{\lambda}$-consequence operator on $\mathscr{L}$ for any cardinal $\lambda$ such that $2\leq\lambda<\kappa$. One might compare this with Example 3.26.
    Theorem 3.39 (Representation Theorem for $r_{\kappa}$-type).
    Let $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S_{r_{\kappa}},\mathcal{P}(\mathscr{L}))$ be a strongly granular semantics for $\mathscr{L}$. Let $S_{r_{\kappa}}=\{(\models_{1},\models_{2})\}$ be such that,
    1. (1)
      $\models_{1}\,\not\subseteq\,\models_{2}$,
    2. (2)
      there exists $\Sigma\cup\Delta\subseteq\mathscr{L}$ such that $\Sigma\subseteq W_{\mathfrak{S}}(\Delta)\cup\Delta$ but $W_{\mathfrak{S}}(\Sigma)\not\subseteq W_{\mathfrak{S}}(\Delta)$,
    3. (3)
      for all $\Lambda\subseteq\mathscr{L}$ with $|\Lambda|\geq\kappa$ there exists $\lambda\in\Lambda$ such that $\pi_{1}(\models_{1}\cap\,\mathbf{M}\times\{\Lambda\})\subseteq\,\pi_{1}(\models_{2}\cap\,\mathbf{M}\times\{\{\lambda\}\})$.
    Then, $W_{\mathfrak{S}}$ is an $r_{\kappa}$-consequence operator.
    Proof.
    Since $\mathfrak{S}$ is strongly granular, by Theorem 3.1, $W_{\mathfrak{S}}$ is monotonic.
    Since $\models_{1}\,\not\subseteq\,\models_{2}$, there exists $m\in\mathbf{M}$ and $\Gamma\subseteq\mathscr{L}$ such that $m\models_{1}\Gamma$ but $m\not\models_{2}\Gamma$. Since $\mathfrak{S}$ is strongly granular, hence granular the 2nd component, there exists $\beta\in\Gamma$ such that $m_{\Gamma}\not\models_{2}\{\beta\}$. So, as $m_{\Gamma}\models_{1}\Gamma$, $\beta\notin W_{\mathfrak{S}}(\Gamma)$, i.e., $\Gamma\not\subseteq W_{\mathfrak{S}}(\Gamma)$. Therefore, $W_{\mathfrak{S}}$ is non-reflexive.
    Next, let us choose any $\Sigma\cup\Delta\subseteq\mathscr{L}$ satisfying (2). Then, $\Sigma\subseteq W_{\mathfrak{S}}(\Delta)\cup\Delta$ but $W_{\mathfrak{S}}(\Sigma)\not\subseteq W_{\mathfrak{S}}(\Delta)$. So, there exists $\alpha\in\mathscr{L}$ such that $\alpha\in W_{\mathfrak{S}}(\Sigma)$ but $\alpha\notin W_{\mathfrak{S}}(\Delta)$. Thus, there exists $n\in\mathbf{M}$ such that $n\models_{1}\Delta$ and $n\not\models_{2}\{\alpha\}$; but, for all $m\in\mathbf{M}$ either $m\not\models_{1}\Sigma$ or $m\models_{2}\{\alpha\}$. We note that $\Sigma\not\subseteq\Delta$ because otherwise, since $n\models_{1}\Delta$, we would have $n\models_{1}\Sigma$. But then $n\models_{2}\{\alpha\}$ - a contradiction. Now, as $\Sigma\not\subseteq\Delta$, there exists $\sigma\in\Sigma$ such that $\sigma\notin\Delta$. However, since $\Sigma\subseteq W_{\mathfrak{S}}(\Delta)\cup\Delta$, $\sigma\in W_{\mathfrak{S}}(\Delta)$. Hence, $W_{\mathfrak{S}}(\Delta)\not\subseteq\Delta$. Therefore, $W_{\mathfrak{S}}$ satisfies non-quasi-closure as well.
    Finally, let $\Lambda\subseteq\mathscr{L}$ such that $|\Lambda|\geq\kappa$. Then, by (3), there exists $\lambda\in\Lambda$ such that $\pi_{1}(\models_{1}\cap\,(\mathbf{M}\times\{\Lambda\}))\,\subseteq\,\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\lambda\}\}))$. We choose any such $\lambda$ and $m\in\mathbf{M}$ such that $m\models_{1}\Lambda$. Then, since $\pi_{1}(\models_{1}\cap\,(\mathbf{M}\times\{\Lambda\}))\,\subseteq\,\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\lambda\}\}))$, $m\models_{2}\{\lambda\}$. Thus, $\lambda\in\Lambda\cap W_{\mathfrak{S}}(\Lambda)$ and hence $\Lambda\not\subseteq\mathscr{L}\setminus W_{\mathfrak{S}}(\Lambda)$. Therefore, $W_{\mathfrak{S}}$ satisfies $\text{non-anti-reflexivity}_{\kappa}$ as well.
    Consequently, $W_{\mathfrak{S}}$ is an $r_{\kappa}$-consequence operator.∎
    The functionally 4-valued characterisation, which was mentioned in Theorem 3.2, works here as well. But, in this case, we can also show that the logical structures induced by $r_{\kappa}$-consequence operators are inferentially 4-valued.
    Theorem 3.40 (Inferential Irreducibility Theorem for $r_{\kappa}$-type).
    Every logical structure induced by an $r_{\kappa}$-consequence operator is inferentially 4-valued.
    Proof.
    We first note that every $r_{\kappa}$-consequence operator has a functionally $4$-valued characterisation by Theorem 3.2. Let $(\mathscr{L},W)$ be a logical structure induced by an $r_{\kappa}$-consequence operator. Suppose that it is not inferentially $4$-valued. So, there exists some functionally $2$-valued or $3$-valued semantics for $\mathscr{L}$, say $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S_{r_{\kappa}},\mathcal{P}(\mathscr{L}))$ such that $W=W_{\mathfrak{S}}$.
    If $\mathfrak{S}$ is functionally $2$-valued, then along the same lines as in the proof of Theorem 3.10, we are done. We skip this, and move on to where $\mathfrak{S}$ is a functionally $3$-valued functional semantics. In this case, without loss of generality, we may assume that $\mathbf{M}\subseteq\{0,1,2\}^{\mathscr{L}}$.
    Then, for $i\in\{1,2\}$, there exists $\emptyset\subsetneq D_{i}\subsetneq\{0,1,2\}$ such that $D_{1}\cup D_{2}\subsetneq\{0,1,2\}$ and for all $m\in\mathbf{M}$, $\Gamma\subseteq\mathscr{L}$, $m\models_{i}\Gamma$ iff $m(\Gamma)\subseteq D_{i}$. If $D_{1}\cup D_{2}$ is a singleton set, then using arguments similar to the one in the proof Theorem 3.10, we are done. Without loss of generality, let $D_{1}\cup D_{2}=\{1,2\}$. Now, $D_{1}\subseteq D_{2}$ implies that $\models_{1}\subseteq\models_{2}$, which by Theorem 3.14 implies that $W$ is a $p$-consequence operator. Similarly, $D_{2}\subseteq D_{1}$, implies that $\models_{1}\subseteq\models_{2}$, which by Theorem 3.7 implies that $W$ is a $q$-consequence operator. Therefore, $D_{1}\not\subseteq D_{2}$ and $D_{2}\not\subseteq D_{1}$ as $W$ is neither a $p$- nor $q$- consequence operator.
    Either $D_{1}\cap D_{2}=\emptyset$ or $D_{1}\cap D_{2}\neq\emptyset$. Suppose $D_{1}\cap D_{2}=\emptyset$. Let $\Gamma\subseteq\mathscr{L}$ such that $|\Gamma|\geq\kappa$ and $\alpha\in\Gamma$. Moreover, let $m\in\mathbf{M}$ such that $m\models_{1}\Gamma$, i.e., $m(\Gamma)\subseteq D_{1}$. Then, as $D_{1}\cap D_{2}=\emptyset$, $m(\alpha)\notin D_{2}$, i.e., $m\not\models_{2}\{\alpha\}$. In other words, $\alpha\notin W_{\mathfrak{S}}(\Gamma)$. Thus $\Gamma\subseteq\mathscr{L}\setminus W_{\mathfrak{S}}(\Gamma)$ for all $|\Gamma|\geq\kappa$, which implies that $W$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$. This is a contradiction to the assumption that $W$ is a $r_{\kappa}$-consequence operator. So, $D_{1}\cap D_{2}\neq\emptyset$.
    If $D_{1}\cap D_{2}=\{1,2\}=D_{1}\cup D_{2}$, then $D_{1}=D_{2}$, contrary to our earlier conclusions that $D_{1}\not\subseteq D_{2}$ and $D_{2}\not\subseteq D_{1}$. Therefore, $D_{1}\cap D_{2}\subsetneq\{1,2\}$. Without loss of generality, let us assume that $D_{1}\cap D_{2}=\{1\}$. Now, as $D_{1}\not\subseteq D_{2}$, there exists $a\in D_{1}$ such that $a\notin D_{2}$. Again, without loss of generality, let $a=2$. So, $D_{1}=\{1,2\}$. Then, as $D_{1}\cup D_{2}\subsetneq\{0,1,2\}$, $D_{2}=\{1\}$ - a contradiction, since $D_{2}\not\subseteq D_{1}$. Thus $W\neq W_{\mathfrak{S}}$. Hence $W$ is inferentially 4-valued. ∎
    Corollary 3.41 (Inferential Characterisation of the Class of Monotonic Logical Structures).
    The class of monotonic logical structures is inferentially 4-valued.
    4 Suszko Reduction: A Generalised Perspective
    In this section, we propose a different notion of semantics than the one in Definition 2.1 and show that every monotonic logical structure has an adequate 2-valued semantics of this kind.
    Definition 4.1 ($S$-semantics).
    Given a set $\mathscr{L}$, a Suszko semantics for $\mathscr{L}$ or $S$-semantics for $\mathscr{L}$, say $\mathsf{S}$, is a tuple $(\mathbf{B},R,\models,\mathcal{P}(\mathscr{L}))$ satisfying the following properties.
    1. (i)
      $\mathbf{B}\subseteq\{0,1\}^{\mathscr{L}}$.
    2. (ii)
      $R\subseteq\mathbf{B}\times\mathbf{B}$.
    3. (iii)
      $\models\,\subseteq\mathbf{B}\times\mathcal{P}(\mathscr{L})$.
    Given an $S$-semantics $\mathsf{S}=(\mathbf{B},R,\models,\mathcal{P}(\mathscr{L}))$, the type-I $S$-entailment relation on $\mathscr{L}$, denoted by $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{\mathsf{S}}_{I}$ is defined as follows:
    $\Gamma\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{\mathsf{S}}_{I}\alpha\iff~\text{for all}~(v,w)\in R,~v\models\Gamma~\text{implies that}~w\models\{\alpha\}$
    The corresponding consequence operator is denoted by $W^{\mathsf{S}}_{I}$. The type-II $S$-entailment relation on $\mathscr{L}$, denoted by $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{\mathsf{S}}_{II}$, is defined as follows.
    $\Gamma\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{\mathsf{S}}_{II}\alpha\iff$ $\text{there exists}~U\subseteq R~\text{with}~(\chi_{\Gamma},\chi_{W(\Gamma)})\in U~\text{such that}$
    $\text{for all}~(v,w)\in U,v\models\Gamma~\text{implies that}~w\models\{\alpha\}$
    The corresponding consequence operator is denoted by $W^{\mathsf{S}}_{II}$.
    Given a logical structure $(\mathscr{L},W)$ and an $S$-semantics $\mathsf{S}$ for $\mathscr{L}$, $(\mathscr{L},W)$ is said to have an adequate type-I $S$-semantics (respectively, adequate type-II $S$-semantics) if $W=W^{\mathsf{S}}_{I}$ (respectively, if $W=W^{\mathsf{S}}_{II}$).
    Remark 4.2.
    The name ‘Suszko semantics’ has already been used in [15, p. 384]. We quote the relevant passage below. Our definition, however, is more general than the one given below.
[…]
    Remark 4.3.
    Let $\mathsf{S}=(\mathbf{B},R,\models,\mathcal{P}(\mathscr{L}))$ an $S$-semantics for a set $\mathscr{L}$ such that $\mathbf{B}=\emptyset$. If there exists $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$ such that $\alpha\notin W^{\mathsf{S}}(\Gamma)$ then there exists $(v,w)\in R$ such that $v\models\Gamma$ but $w\not\models\{\alpha\}$. This, however, is a contradiction since $R=\emptyset$ as $\emptyset\subseteq R\subseteq\mathbf{B}\times\mathbf{B}=\emptyset\times\emptyset=\emptyset$. Hence, $W_{\mathfrak{S}}(\Gamma)=\mathscr{L}$ for all $\Gamma\subseteq\mathscr{L}$.
    The following theorem shows that every logical structure induced by a semantics (in the sense of Definition 2.1) has an adequate type-I $S$-semantics.
    Theorem 4.4 (Suszko Reduction for Logical Structures Induced by a Semantics).
    If a logical structure has an adequate semantics, it has an adequate type-I $S$-semantics as well.
    Proof.
    Let $(\mathscr{L},W)$ be a logical structure that is adequate with respect to a semantics $\mathfrak{S}=(\textbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))$ for $\mathscr{L}$. So, $W=W_{\mathfrak{S}}$.
    If $\mathbf{M}=\emptyset$ then we consider any $\mathsf{S}$-semantics $\mathsf{S}=(\mathbf{B},R,\models,\mathcal{P}(\mathscr{L}))$ for $\mathscr{L}$ such that $\mathbf{B}=\emptyset$. Then by, Remark 2.2 and Remark 4.3 $W_{\mathfrak{S}}(\Gamma)=W^{\mathsf{S}}(\Gamma)=\mathscr{L}$ for all $\Gamma\subseteq\mathscr{L}$. Therefore, in this case, $W_{\mathfrak{S}}=W^{\mathsf{S}}$.
    If, on the other hand, $\mathbf{M}\neq\emptyset$, then for each $(m,i)\in\mathbf{M}\times I$, we choose $v^{m}_{i}\in\{0,1\}^{\mathscr{L}}$ and define a $S$-semantics for $\mathscr{L}$, $\mathsf{S}^{\mathfrak{S}}=(\textbf{B}^{\mathfrak{S}},R^{\mathfrak{S}},\models,\mathcal{P}(\mathscr{L}))$, as follows.
    1. (1)
      $\mathbf{B}^{\mathfrak{S}}=\{v^{m}_{i}\in\{0,1\}^{\mathscr{L}}:(m,i)\in\mathbf{M}\times I\}$.
    2. (2)
      $R^{\mathfrak{S}}=\{(v^{m}_{i},v^{m}_{j}):(\models_{i},\models_{j})\in S~\text{and}~m\in\mathbf{M}\}$.
    3. (3)
      For all $(m,i)\in\mathbf{M}\times I$, $v^{m}_{i}\models\Gamma$ iff $m\models_{i}\Gamma$.
    We now show that $W_{\mathfrak{S}}=W^{\mathsf{S}^{\mathfrak{S}}}_{I}$.
    Let $\Gamma\subseteq\mathscr{L}$ such that $W_{\mathfrak{S}}(\Gamma)\not\subseteq W^{\mathsf{S}^{\mathfrak{S}}}_{I}(\Gamma)$. Then there exists $\alpha\in W_{\mathfrak{S}}(\Gamma)$ such that $\alpha\notin W^{\mathsf{S}^{\mathfrak{S}}}_{I}(\Gamma)$. So, there exists $(v^{m}_{i},v^{m}_{j})\in R^{\mathfrak{S}}$ such that $v^{m}_{i}\models\Gamma$, but $v^{m}_{j}\not\models\{\alpha\}$. Now, $v^{m}_{i}\models\Gamma$ implies that $m\models_{i}\Gamma$ and $v^{m}_{j}\not\models\{\alpha\}$ implies that $m\not\models_{j}\{\alpha\}$. In other words, $\alpha\notin W_{\mathfrak{S}}(\Gamma)$ - a contradiction. Thus $W_{\mathfrak{S}}(\Gamma)\subseteq W^{\mathsf{S}^{\mathfrak{S}}}_{I}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$.
    Next, let $\Gamma\subseteq\mathscr{L}$ such that $W^{\mathsf{S}^{\mathfrak{S}}}_{I}(\Gamma)\not\subseteq W_{\mathfrak{S}}(\Gamma)$, Then there exists $\alpha\in W^{\mathsf{S}^{\mathfrak{S}}}_{I}(\Gamma)$ such that $\alpha\notin W_{\mathfrak{S}}(\Gamma)$. So, there exists $m\in\mathbf{M}$ and $(\models_{i},\models_{j})\in S$ such that $m\models_{i}\Gamma$, but $m\not\models_{j}\{\alpha\}$. Now, $m\models_{i}\Gamma$ implies that $v^{m}_{i}\models\Gamma$, and $m\not\models_{j}\{\alpha\}$ implies that $v^{m}_{j}\not\models\{\alpha\}$. In other words, $\alpha\notin W^{\mathsf{S}^{\mathfrak{S}}}_{I}(\Gamma)$ - a contradiction. Thus $W^{\mathsf{S}^{\mathfrak{S}}}_{I}(\Gamma)\subseteq W_{\mathfrak{S}}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$. Therefore, $W_{\mathfrak{S}}=W^{\mathsf{S}^{\mathfrak{S}}}_{I}$. ∎
    The converse of the above theorem is established below.
    Theorem 4.5 (Semantic Counterpart of an $S$-semantics).
    If a logical structure has an adequate type-I $S$-semantics, then it has an adequate semantics as well.
    Proof.
    Let $(\mathscr{L},W)$ be a logical structure that has an adequate type-I $S$-semantics $\mathsf{S}=(\textbf{B},R,\models,\mathcal{P}(\mathscr{L}))$ for $\mathscr{L}$. So, $W=W^{\mathsf{S}}_{I}$.
    If $\mathbf{B}=\emptyset$, then we consider any semantics $\mathfrak{S}=(\mathbf{M},\{\models_{i}\}_{i\in I},S,\mathcal{P}(\mathscr{L}))$ for $\mathscr{L}$ such that $\mathbf{M}=\emptyset$. Then by, Remark 4.3 and Remark 2.2 $W^{\mathsf{S}}_{I}(\Gamma)=\mathscr{L}=W_{\mathfrak{S}}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$. Therefore, in this case, $W_{\mathfrak{S}}=W^{\mathsf{S}}_{I}$.
    If, on the other hand, $\mathbf{B}\neq\emptyset$, then we define a semantics for $\mathscr{L}$, $\mathfrak{S}_{\mathsf{S}}=(\textbf{M}_{\mathsf{S}},\{\models_{v}\}_{v\in\mathbf{B}},S_{\mathsf{S}},\mathcal{P}(\mathscr{L}))$ as follows.
    1. (1)
      $\mathbf{M}_{\mathsf{S}}=\{\ast\}$ where $\ast\notin\mathbf{B}$.
    2. (2)
      For each $v\in\mathbf{B}$, $\ast\models_{v}\Gamma~\text{iff}~v\models\Gamma$.
    3. (3)
      $S_{\mathsf{S}}=\{(\models_{v},\models_{w}):(v,w)\in R\}$.
    We now claim that $W_{\mathfrak{S}_{\mathsf{S}}}=W^{\mathsf{S}}_{I}$.
    Let $\Gamma\subseteq\mathscr{L}$ such that $W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)\not\subseteq W^{\mathsf{S}}_{I}(\Gamma)$. Then there exists $\alpha\in W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)$ such that $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$. So, there exists $(v,w)\in R$ such that $v\models\Gamma$, but $w\not\models\{\alpha\}$. Now, $v\models\Gamma$ implies that $\ast\models_{v}\Gamma$ and $w\not\models\{\alpha\}$ implies that $\ast\not\models_{w}\{\alpha\}$. In other words, $\alpha\notin W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)$ - a contradiction. Thus $W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)\subseteq W^{\mathsf{S}}_{I}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$.
    Next, let $\Gamma\subseteq\mathscr{L}$ such that $W^{\mathsf{S}}_{I}(\Gamma)\not\subseteq W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)$, Then there exists $\alpha\in W^{\mathsf{S}}_{I}(\Gamma)$ such that $\alpha\notin W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)$. So, there exists $(\models_{v},\models_{w})\in S$ such that $\ast\models_{v}\Gamma$, but $\ast\not\models_{w}\{\alpha\}$. Now, $\ast\models_{v}\Gamma$ implies that $v\models\Gamma$, and $\ast\not\models_{w}\{\alpha\}$ implies that $w\not\models\{\alpha\}$. In other words, $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$ - a contradiction. Thus $W^{\mathsf{S}}_{I}(\Gamma)\subseteq W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$. Therefore, $W_{\mathfrak{S}_{\mathsf{S}}}=W^{\mathsf{S}}_{I}$. ∎
    4.1 Suszko Reduction for Monotonic Logical Structures
    Definition 4.6 (Atomic $S$-semantics, Bivalent Logical Structure).
    Given a set $\mathscr{L}$ and an $S$-semantics $\mathsf{S}=(\mathbf{B},R,\models,\mathcal{P}(\mathscr{L}))$ for $\mathscr{L}$, it is said to be atomic if for all $\Gamma\subseteq\mathscr{L}$ and $v\in\mathbf{B}$, $v\models\Gamma$ iff $v\models\{\beta\}$ for all $\beta\in\Gamma$. We call a logical structure bivalent if it is adequate with respect to type-I (or type II) $S$-semantics.
    Remark 4.7.
    One may compare the notion of atomic $S$-semantics with that of granular semantics defined in Definition 2.14.
    Given a set $\mathscr{L}$ and an atomic $S$-semantics $\mathsf{S}=(\mathbf{B},R,\models,\mathcal{P}(\mathscr{L}))$ for $\mathscr{L}$, one can construct the following $S$-semantics $\mathsf{S}_{2}=(\mathbf{B}_{2},R_{2},\models_{2},\mathcal{P}(\mathscr{L}))$ for $\mathscr{L}$ such that the following conditions hold.
    1. $\bullet$
      $\mathbf{B}_{2}=\{\nu_{v}:v\in\mathbf{B}\}$ where $\nu_{v}:\mathscr{L}\to\{0,1\}$ is defined as follows.
      $\nu_{v}(\beta)=\begin{cases}1&\text{if}~v\models\{\beta\}\\ 0&\text{otherwise}\end{cases}$
    2. $\bullet$
      $R_{2}=\{(\nu_{v},\nu_{w}):(v,w)\in R\}$.
    3. $\bullet$
      For all $\nu_{v}\in\mathbf{B}_{2}$, $\nu_{v}\models_{2}\Gamma$ iff $\nu_{v}(\Gamma)\subseteq\{1\}$
    It is easy to show that $W^{\mathsf{S}}=W^{\mathsf{S}_{2}}$. Due to this, we will henceforth consider only those atomic $S$-semantics for which $v\models\Gamma$ iff $v(\Gamma)\subseteq\{1\}$. This gives rise to the following definition.
    Definition 4.8 (Normal $S$-semantics).
    An atomic $S$-semantics $\mathsf{S}=(\mathbf{B},R,\models,\mathcal{P}(\mathscr{L}))$ for a set $\mathscr{L}$ is said to be normal if for all $v\in\mathbf{B}$ and $\Gamma\subseteq\mathscr{L}$, $v\models\Gamma$ iff $v(\Gamma)\subseteq\{1\}$
    In the following two theorems, we show that every monotonic logical structure is bivalent.
    Theorem 4.9 (Suszko Reduction for Monotonic Logical Structures (Part I)).
    Let $\mathscr{L}$ be a set and $\mathsf{S}=(\mathbf{B},R,\models,\mathcal{P}(\mathscr{L}))$ a normal $S$-semantics for $\mathscr{L}$. Then $W^{\mathsf{S}}_{I}$ is monotonic.
    Proof.
    Suppose the contrary. Let $\Gamma\subseteq\Sigma\subseteq\mathscr{L}$ such that $W^{\mathsf{S}}_{I}(\Gamma)\not\subseteq W^{\mathsf{S}}_{I}(\Sigma)$. So, there exists $\alpha\in W^{\mathsf{S}}_{I}(\Gamma)$ such that $\alpha\notin W^{\mathsf{S}}_{I}(\Sigma)$. Then, there exists $(v,w)\in R$ such that $v\models\Sigma$ but $w\not\models\{\alpha\}$, i.e., $v(\Sigma)\subseteq\{1\}$ but $w(\alpha)\neq 1$. Now, as $\Gamma\subseteq\Sigma$, $v(\Sigma)\subseteq\{1\}$ implies that $v(\Gamma)\subseteq\{1\}$. Then, since $w(\alpha)\neq 1$, $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$ - a contradiction. Hence, $W^{\mathsf{S}}_{I}$ is monotonic. ∎
    Theorem 4.10 (Suszko Reduction for Monotonic Logical Structures (Part II)).
    Let $(\mathscr{L},W)$ be a monotonic logical structure and $\mathsf{S}=(\mathbf{B},R_{m},\models,\mathcal{P}(\mathscr{L}))$ be a normal $S$-semantics for $\mathscr{L}$ where
    $\mathbf{B}=\{\chi_{\Gamma}:\Gamma\subseteq\mathscr{L}\}$
    and,
    $R_{m}=\{(\chi_{\Gamma},\chi_{W(\Gamma)}):\Gamma\subseteq\mathscr{L}\}$
    Then, $W=W^{\mathsf{S}}_{I}$. Here $\chi_{\Gamma}:\mathscr{L}\to\{0,1\}$ defined as follows.
    $\chi_{\Gamma}(\beta)=\begin{cases}1&\text{if}~\beta\in\Gamma\\ 0&\text{otherwise}\end{cases}$
    Proof.
    Let $\Gamma\subseteq\mathscr{L}$ such that $W(\Gamma)\not\subseteq W^{\mathsf{S}}_{I}(\Gamma)$. Then, there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$. So, there exists $(\chi_{\Sigma},\chi_{W(\Sigma)})\in R_{m}$ such that $\chi_{\Sigma}\models\Gamma$ but $\chi_{W(\Sigma)}\not\models\{\alpha\}$, i.e., $\chi_{\Sigma}(\Gamma)\subseteq\{1\}$ but $\chi_{W(\Sigma)}(\alpha)\neq 1$. Now, $\chi_{\Sigma}(\Gamma)\subseteq\{1\}$ implies that $\Gamma\subseteq\Sigma$ and $\chi_{W(\Sigma)}(\alpha)\neq 1$ implies that $\alpha\notin W(\Sigma)$. But, since $\Gamma\subseteq\Sigma$, and $W$ is monotonic, $W(\Gamma)\subseteq W(\Sigma)$. Then, as $\alpha\in W(\Gamma)$, $\alpha\in W(\Sigma)$ - a contradiction. Thus $W(\Gamma)\subseteq W^{\mathsf{S}}_{I}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$.
    Next, let $\Gamma\subseteq\mathscr{L}$ such that $W^{\mathsf{S}}_{I}(\Gamma)\not\subseteq W(\Gamma)$. Then, there exists $\alpha\in W^{\mathsf{S}}_{I}(\Gamma)$ such that $\alpha\notin W(\Gamma)$. Now $(\chi_{\Gamma},\chi_{W(\Gamma)})\in R_{m}$, $\chi_{\Gamma}(\Gamma)\subseteq\{1\}$ but $\chi_{W(\Gamma)}(\alpha)\neq 1$. So, $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$ - a contradiction. Consequently, $W^{\mathsf{S}}_{I}(\Gamma)\subseteq W(\Sigma)$ for all $\Gamma\subseteq\mathscr{L}$. Hence $W=W^{\mathsf{S}}_{I}$. ∎
    Since every monotonic logical structure has an adequate type-I normal $S$-semantics, any $q$-, $p$-, $s_{\kappa}$- or $r_{\kappa}$-type logical structure has an adequate type-I normal $S$-semantics. For $q$-, $p$- and $s_{\kappa}$-type logical structures, however, the elements of $\mathbf{B}$ and $R$ satisfies some additional properties, as demonstrated in the following theorems.
    Theorem 4.11 (Suszko Reduction for $q$-type (Part I)).
    Let $\mathscr{L}$ be a set and $\mathsf{S}=(\mathbf{B},R,\models,\mathcal{P}(\mathscr{L}))$ be a normal $S$-semantics such that,
    $(v,w)\in R~\text{iff for all}~\alpha\in\mathscr{L},w(\alpha)=1~\text{implies that}~v(\alpha)=1$
    Then, $W^{\mathsf{S}}_{I}$ is a $q$-consequence operator.
    Proof.
    Let $\Gamma\cup\Sigma\subseteq\mathscr{L}$ such that $\Gamma\subseteq W^{\mathsf{S}}_{I}(\Sigma)\cup\Sigma$ but $W^{\mathsf{S}}_{I}(\Gamma)\not\subseteq W^{\mathsf{S}}_{I}(\Sigma)$. So, there exists $\alpha\in W^{\mathsf{S}}_{I}(\Gamma)$ such that $\alpha\notin W^{\mathsf{S}}_{I}(\Sigma)$. Then, there exists $(v,w)\in R$ such that $v(\Sigma)\subseteq\{1\}$ but $w(\alpha)\neq 1$.
    Let $\beta\in W^{\mathsf{S}}_{I}(\Sigma)$. Then, as $v(\Sigma)\subseteq\{1\}$ and $(v,w)\in R$, $w(\beta)=1$. Now, by the definition of $R$, since $(v,w)\in R$, and $w(\beta)=1$, it follows that $v(\beta)=1$ as well. Consequently, $v(W^{\mathsf{S}}_{I}(\Sigma)\cup\Sigma)\subseteq\{1\}$. Since $\Gamma\subseteq W^{\mathsf{S}}_{I}(\Sigma)\cup\Sigma$, $v(\Gamma)\subseteq\{1\}$ as well. Now, as $w(\alpha)\neq 1$, $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$ - a contradiction. Thus $W^{\mathsf{S}}_{I}(\Gamma)\subseteq W^{\mathsf{S}}_{I}(\Sigma)$. Hence for all $\Gamma\cup\Sigma\subseteq\mathscr{L}$, $\Gamma\subseteq W(\Sigma)\cup\Sigma$ implies that $W(\Gamma)\subseteq W(\Sigma)$. So, by Theorem 3.6(5), $W^{\mathsf{S}}_{I}$ is a $q$-consequence operator. ∎
    Theorem 4.12 (Suszko Reduction for $q$-type (Part II)).
    Let $(\mathscr{L},W)$ be $q$-type logical structure and $\mathsf{S}=(\mathbf{B},R^{q},\models,\mathcal{P}(\mathscr{L}))$ be a normal $S$-semantics for $\mathscr{L}$, where
    $\mathbf{B}=\{\chi_{W(\Gamma)\cup\Gamma}:\Gamma\subseteq\mathscr{L}\}\cup\{\chi_{W(\Gamma)}:\Gamma\subseteq\mathscr{L}\}$
    and,
    $R^{q}=\{(\chi_{W(\Gamma)\cup\Gamma},\chi_{W(\Gamma)}):\Gamma\subseteq\mathscr{L}\}$
    Then, $W=W^{\mathsf{S}}_{I}$. $\chi_{\Gamma}$’s are defined as in Theorem 4.10.
    Proof.
    Suppose $\Gamma\subseteq\mathscr{L}$ such that $W(\Gamma)\not\subseteq W^{\mathsf{S}}_{I}(\Gamma)$. So, $\alpha\in W(\Gamma)$ such that $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$. Now, as $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$, there exists $(\chi_{W(\Sigma)\cup\Sigma},\chi_{W(\Sigma)})\in R^{q}$ such that $\chi_{W(\Sigma)\cup\Sigma}(\Gamma)\subseteq\{1\}$ but $\chi_{W(\Sigma)}(\alpha)\neq 1$. Now, $\chi_{W(\Sigma)\cup\Sigma}(\Gamma)\subseteq\{1\}$ implies that $\Gamma\subseteq W(\Sigma)\cup\Sigma$, while $\chi_{W(\Sigma)}(\alpha)\neq 1$ implies that $\alpha\notin W(\Sigma)$. Since $\Gamma\subseteq W(\Sigma)\cup\Sigma$ and $W$ is a $q$-consequence operator, by Theorem 3.6(5), $W(\Gamma)\subseteq W(\Sigma)$. Then, as $\alpha\in W(\Gamma)$, $\alpha\in W(\Sigma)$ - a contradiction. Thus, $W(\Gamma)\subseteq W^{\mathsf{S}}_{I}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$.
    Next, let $\Gamma\subseteq\mathscr{L}$ and $\alpha\in W^{\mathsf{S}}_{I}(\Gamma)$. Then, $(\chi_{W(\Gamma)\cup\Gamma},\chi_{W(\Gamma)})\in R^{q}$ and $\chi_{W(\Gamma)\cup\Gamma}(\Gamma)\subseteq\{1\}$. Since $\alpha\in W^{\mathsf{S}}_{I}(\Gamma)$, $\chi_{W(\Gamma)}(\alpha)=1$, i.e., $\alpha\in W(\Gamma)$. Thus $W^{\mathsf{S}}_{I}(\Gamma)\subseteq W(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$. Hence $W=W^{\mathsf{S}}_{I}$. ∎
    Theorem 4.13 (Suszko Reduction for $p$-type (Part I)).
    Let $\mathscr{L}$ be a set and $\mathsf{S}=(\mathbf{B},R,\models,\mathcal{P}(\mathscr{L}))$ a normal $S$-semantics such that,
    $(v,w)\in R~\text{iff for all}~\alpha\in\mathscr{L},v(\alpha)=1~\text{implies that}~w(\alpha)=1$
    Then, $W^{\mathsf{S}}_{I}$ is a $p$-consequence operator.
    Proof.
    Let there exists $\Gamma\cup\Sigma\subseteq\mathscr{L}$ such that $\Gamma\subseteq\Sigma$ but $\Gamma\cup W^{\mathsf{S}}_{I}(\Gamma)\not\subseteq W^{\mathsf{S}}_{I}(\Sigma)$. So, there exists $\alpha\in\Gamma\cup W^{\mathsf{S}}_{I}(\Gamma)$ such that $\alpha\notin W^{\mathsf{S}}_{I}(\Sigma)$. Then, there exists $(v,w)\in R$ such that $v(\Sigma)\subseteq\{1\}$ but $w(\alpha)\neq 1$.
    Since $\Gamma\subseteq\Sigma$ and $v(\Sigma)\subseteq\{1\}$, $v(\Gamma)\subseteq\{1\}$ as well. Also, since $w(\alpha)\neq 1$, $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$. So, $\alpha\in\Gamma$. Then $v(\alpha)=1$, since $v(\Gamma)\subseteq\{1\}$. Now, by the definition of $R$, since $(v,w)\in R$, $v(\alpha)=1$ implies that $w(\alpha)=1$ - a contradiction. Hence for all $\Gamma\cup\Sigma\subseteq\mathscr{L}$, $\Gamma\subseteq\Sigma$ implies that $\Gamma\cup W_{I}^{\mathsf{S}}(\Gamma)\subseteq W_{I}^{\mathsf{S}}(\Sigma)$. So, by Theorem 3.13, $W^{\mathsf{S}}_{I}$ is a $p$-consequence operator. ∎
    Theorem 4.14 (Suszko Reduction for $p$-type (Part II)).
    Let $(\mathscr{L},W)$ be a $p$-type logical structure and $\mathsf{S}=(\mathbf{B},R^{p},\models,\mathcal{P}(\mathscr{L}))$ a normal $S$-semantics for $\mathscr{L}$, where
    $\mathbf{B}=\{\chi_{\Gamma}:\Gamma\subseteq\mathscr{L}\}\cup\{\chi_{\Gamma\cup W(\Gamma)}:\Gamma\subseteq\mathscr{L}\}$
    and
    $R^{p}=\{(\chi_{\Gamma},\chi_{\Gamma\cup W(\Gamma)}):\Gamma\subseteq\mathscr{L}\}$
    Then, $W=W^{\mathsf{S}}_{I}$.
    Proof.
    Let $\Gamma\subseteq\mathscr{L}$ such that $W(\Gamma)\not\subseteq W^{\mathsf{S}}_{I}(\Gamma)$. Then, there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$. So, there exists $(\chi_{\Sigma},\chi_{\Sigma\cup W(\Sigma)})\in R^{p}$ such that $\chi_{\Sigma}(\Gamma)\subseteq\{1\}$ but $\chi_{\Sigma\cup W(\Sigma)}(\alpha)\neq 1$. But, $\chi_{\Sigma}(\Gamma)\subseteq\{1\}$ implies that $\Gamma\subseteq\Sigma$, while $\chi_{\Sigma\cup W(\Sigma)}(\alpha)\neq 1$ implies that $\alpha\notin\Sigma\cup W(\Sigma)$, and so $\alpha\notin W(\Sigma)$. Since $\Gamma\subseteq\Sigma$, and $W$ is a $p$-consequence operator, by Theorem 3.13, $\Gamma\cup W(\Gamma)\subseteq W(\Sigma)$. Then, as $\alpha\notin W(\Sigma)$, $\alpha\notin\Gamma\cup W(\Gamma)$, which implies that $\alpha\notin W(\Gamma)$ - a contradiction. Thus, $W(\Gamma)\subseteq W^{\mathsf{S}}_{I}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$.
    Next, let $\Gamma\subseteq\mathscr{L}$ such that $W^{\mathsf{S}}_{I}(\Gamma)\not\subseteq W(\Gamma)$. Then, there exists $\alpha\in W^{\mathsf{S}}_{I}(\Gamma)$ such that $\alpha\notin W(\Gamma)$. Since $W$ is a $p$-consequence operator, and hence reflexive, $\alpha\notin W(\Gamma)$ implies that $\alpha\notin\Gamma$ as well. Thus, $\alpha\notin\Gamma\cup W(\Gamma)$.
    Now, $(\chi_{\Gamma},\chi_{\Gamma\cup W(\Gamma)})\in R^{p}$, $\chi_{\Gamma}(\Gamma)\subseteq\{1\}$ but $\chi_{\Gamma\cup W(\Gamma)}(\alpha)\neq 1$. Hence, $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$ - a contradiction. Thus, $W^{\mathsf{S}}_{I}(\Gamma)\subseteq W(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$. Hence $W=W^{\mathsf{S}}_{I}$. ∎
    Theorem 4.15 (Suszko Reduction for $s^{\mathsf{K}}_{\kappa}$-type (Part I)).
    Let $\mathscr{L}$ be a set and $\mathsf{K}\subseteq\mathcal{P}(\mathscr{L})$ be internally $\kappa$ for some cardinal $\kappa$. Let $\mathsf{S}=(\mathbf{B},R,\models,\mathcal{P}(\mathscr{L}))$ be a normal $S$-semantics such that for each $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$ with $\alpha\in\Gamma$ and $\Gamma\in\mathsf{K}$, there exists $(v,w)\in R$ such that $v(\Gamma)\subseteq\{1\}$ but $w(\alpha)\notin\{1\}$. Then $W^{\mathsf{S}}_{I}$ is an $s^{\mathsf{K}}_{\kappa}$-consequence operator.
    Proof.
    We first note that $W^{\mathsf{S}}_{I}$ is monotonic, by Theorem 4.9.
    Now, let $\Gamma\in\mathsf{K}$, $\alpha\in\Gamma$ and $(v,w)\in R$ such that $v(\Gamma)\subseteq\{1\}$ and $w(\alpha)\neq 1$ (such a $(v,w)$ exists by hypothesis). Thus, $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$, which implies that $\Gamma\subseteq\mathscr{L}\setminus W^{\mathsf{S}}_{I}(\Gamma)$. Hence, $\Gamma\subseteq\mathscr{L}\setminus W^{\mathsf{S}}_{I}(\Gamma)$ for all $\Gamma\in\mathsf{K}$. Thus, $W^{\mathsf{S}}_{I}$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$ and hence an $s^{\mathsf{K}}_{\kappa}$-consequence operator. ∎
    Theorem 4.16 (Suszko Reduction for $s^{\mathsf{K}}_{\kappa}$-type (Part II)).
    Let $(\mathscr{L},W)$ be an $s^{\mathsf{K}}_{\kappa}$-type logical structure and $\mathsf{S}=(\mathbf{B},R_{s^{\mathsf{K}}_{\kappa}},\models,\mathcal{P}(\mathscr{L}))$ a normal $S$-semantics for $\mathscr{L}$ satisfying the following conditions.
    1. $\bullet$
      $\mathbf{B}=\mathbf{B}_{1}\cup\mathbf{B}_{2}\cup\mathbf{B}_{3}$, where
      $\mathbf{B}_{1}=\{\chi_{\mathscr{L}\setminus W(\Gamma)}:\Gamma\in\mathsf{K}~\text{and}~\mathscr{L}\setminus W(\Gamma)\in\mathsf{K}\}$
      $\mathbf{B}_{2}=\{\chi_{W(\Gamma)}:\Gamma\in\mathsf{K}~\text{and}~\mathscr{L}\setminus W(\Gamma)\in\mathsf{K}\}$
      $\mathbf{B}_{3}=\{\chi_{\Gamma}:\Gamma\notin\mathsf{K}~\text{or}~\mathscr{L}\setminus W(\Gamma)\notin\mathsf{K}\}$
    2. $\bullet$
      $R_{s^{\mathsf{K}}_{\kappa}}=R_{1}\cup R_{2}$, where
      $R_{1}=\{(\chi_{\mathscr{L}\setminus W(\Gamma)},\chi_{W(\Gamma)}):\Gamma\in\mathsf{K}~\text{and}~\mathscr{L}\setminus W(\Gamma)\in\mathsf{K}\}$
      $R_{2}=\{(\chi_{\Gamma},\chi_{W(\Gamma)}):\Gamma\notin\mathsf{K}~\text{or}~\mathscr{L}\setminus W(\Gamma)\notin\mathsf{K}\}$
    Then, $W=W^{\mathsf{S}}_{I}$.
    Proof.
    Let for some $\Gamma\subseteq\mathscr{L}$, $W(\Gamma)\not\subseteq W^{\mathsf{S}}_{I}(\Gamma)$. So, there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$. Then, the following cases arise.
    1. (a)
      There exists $(\chi_{\mathscr{L}\setminus W(\Sigma)},\chi_{W(\Sigma)})\in R_{1}$ such that $\chi_{\mathscr{L}\setminus W(\Sigma)}(\Gamma)\subseteq\{1\}$ but $\chi_{W(\Sigma)}(\alpha)\neq 1$.
    2. (b)
      There exists $(\chi_{\Sigma},\chi_{W(\Sigma)})\in R_{2}$ such that $\chi_{\Sigma}(\Gamma)\subseteq\{1\}$ but $\chi_{W(\Sigma)}(\alpha)\neq 1$.
    We discuss each case one by one.
    Case 4.16.1:
    $\chi_{\mathscr{L}\setminus W(\Sigma)}(\Gamma)\subseteq\{1\}$ implies that $\Gamma\subseteq\mathscr{L}\setminus W(\Sigma)$, while $\chi_{W(\Sigma)}(\alpha)\neq 1$ implies that $\alpha\notin W(\Sigma)$. Now,
    $\Gamma\subseteq\mathscr{L}\setminus W(\Sigma)$ $\implies W(\Gamma)\subseteq W(\mathscr{L}\setminus W(\Sigma))$ $\text{(by monotonicity of}~W)$
    $\implies W(\Gamma)\subseteq W(\Sigma)$ (by Theorem 3.23(1), since $\mathscr{L}\setminus W(\Sigma)\in\mathsf{K})$
    $\implies\alpha\in W(\Sigma)$ $\text{(since}~\alpha\in W(\Gamma))$
    This is a contradiction.
    Case 4.16.2:
    Since, $\chi_{\Sigma}(\Gamma)\subseteq\{1\}$, $\Gamma\subseteq\Sigma$. Then, by Theorem 4.9, since $W$ is monotone, $W(\Gamma)\subseteq W(\Sigma)$. So, $\alpha\in W(\Gamma)$ implies that $\alpha\in W(\Sigma)$, i.e. $\chi_{W(\Sigma)}(\alpha)\neq 1$ - a contradiction.
    Therefore, $W(\Gamma)\subseteq W^{\mathsf{S}}_{I}(\Gamma)$, for all $\Gamma\subseteq\mathscr{L}$.
    Next, let $\Gamma\subseteq\mathscr{L}$ and $\alpha\in W^{\mathsf{S}}_{I}(\Gamma)$. If $\Gamma\in\mathsf{K}$ and $\mathscr{L}\setminus W(\Gamma)\in\mathsf{K}$, then $(\chi_{\mathscr{L}\setminus W(\Gamma)},\chi_{W(\Gamma)})\in R_{s^{\mathsf{K}}_{\kappa}}$. Since $W$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$, and $\Gamma\in\mathsf{K}$, $\Gamma\subseteq\mathscr{L}\setminus W(\Gamma)$, which implies that $\chi_{\mathscr{L}\setminus W(\Gamma)}(\Gamma)\subseteq\{1\}$. Therefore, since $\alpha\in W^{\mathsf{S}}_{I}(\Gamma)$, we have $\chi_{W(\Gamma)}(\alpha)=1$, i.e., $\alpha\in W(\Gamma)$. If, on the other hand, $\Gamma\notin\mathsf{K}$ or $\mathscr{L}\setminus W(\Gamma)\notin\mathsf{K}$, then $(\chi_{\Gamma},\chi_{W(\Gamma)})\in R_{s^{\mathsf{K}}_{\kappa}}$. Since $\chi_{\Gamma}(\Gamma)\subseteq\{1\}$ and $\alpha\in W^{\mathsf{S}}_{I}(\Gamma)$, $\chi_{W(\Gamma)}(\alpha)=1$, i.e., $\alpha\in W(\Gamma)$.
    Thus, $W^{\mathsf{S}}_{I}(\Gamma)\subseteq W(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$. Hence, $W=W^{\mathsf{S}}_{I}$. ∎
    4.2 Suszko Reduction for Non-monotonic Logical Structures
    We know from Theorem 2.6 and Theorem 4.5 that every logical structure, and hence any non-monotonic logical structure, has an adequate type-I $S$-semantics.
    Theorem 4.17.
    Let $(\mathscr{L},W)$ be a logical structure and $\mathsf{S}=(\mathbf{B},R,\models,\mathcal{P}(\mathscr{L}))$ an atomic $S$-semantics for $\mathscr{L}$ such that $W=W^{\mathsf{S}}_{I}$. Then, for all $\Gamma\subseteq\mathscr{L}$, $W(\Gamma)=\mathscr{L}$ or $\bigcup_{\begin{subarray}{c}\Gamma_{0}\subseteq\Gamma\\ \Gamma_{0}~\text{is finite}\end{subarray}}W(\Gamma_{0})\subseteq W(\Gamma)$.
    Proof.
    Let $\Gamma\subseteq\mathscr{L}$ such that $W(\Gamma)\neq\mathscr{L}$ and $\alpha\notin W(\Gamma)$. Then, since $W=W^{\mathsf{S}}_{I}$, $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$. So, there exists $(v,w)\in R$ such that $v\models\Gamma$ but $w\not\models\{\alpha\}$. Since $\mathsf{S}$ is atomic, $v\models\Gamma$ implies that $v\models\Gamma_{0}$ for all finite $\Gamma_{0}\subseteq\Gamma$. Now, as $w\not\models\{\alpha\}$, $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma_{0})$ for all finite $\Gamma_{0}\subseteq\Gamma$, i.e., $\alpha\notin\displaystyle\bigcup_{\begin{subarray}{c}\Gamma_{0}\subseteq\Gamma\\ \Gamma_{0}~\text{is finite}\end{subarray}}W^{\mathsf{S}}_{I}(\Gamma_{0})$. Since $W=W^{\mathsf{S}}_{I}$, this implies that $\alpha\notin\displaystyle\bigcup_{\begin{subarray}{c}\Gamma_{0}\subseteq\Gamma\\ \Gamma_{0}~\text{is finite}\end{subarray}}W(\Gamma_{0})$. Thus, $\bigcup_{\begin{subarray}{c}\Gamma_{0}\subseteq\Gamma\\ \Gamma_{0}~\text{is finite}\end{subarray}}W(\Gamma_{0})\subseteq W(\Gamma)$. Hence, for all $\Gamma\subseteq\mathscr{L}$, $W(\Gamma)=\mathscr{L}$ or $\bigcup_{\begin{subarray}{c}\Gamma_{0}\subseteq\Gamma\\ \Gamma_{0}~\text{is finite}\end{subarray}}W(\Gamma_{0})\subseteq W(\Gamma)$. ∎
    However, there exists non-monotonic logical structure $(\mathscr{L},W)$ such that $W$ does not satisfy the property mentioned in the above theorem. The following gives an example of such a logical structure.
    Example 4.18.
    Let $(\mathscr{L},W)$ be a logical structure where $\mathscr{L}=\mathbb{Q}$ (the set of rational numbers) and $W:\mathcal{P}(\mathscr{L})\to\mathcal{P}(\mathscr{L})$ is defined as follows.
    $W(\Gamma)=\begin{cases}\mathbb{N}&\text{if}~\Gamma~\text{is finite}\\ \mathscr{L}\setminus\Gamma&\text{if}~\Gamma~\text{is infinite and}~\Gamma\subsetneq\mathscr{L}\\ \emptyset&\text{otherwise (i.e., if}~\Gamma=\mathscr{L})\end{cases}$
    Clearly, $W(\Gamma)\neq\mathscr{L}$ for all $\Gamma\subseteq\mathscr{L}$. Choose $\mathbb{Q}$ and note that,
    $\bigcup_{\begin{subarray}{c}\Gamma_{0}\subseteq\mathbb{Q}\\ \Gamma_{0}~\text{is finite}\end{subarray}}W(\Gamma_{0})=\mathbb{N}\not\subseteq\emptyset=W(\mathbb{Q})$
    However, $(\mathscr{L},W)$ is non-monotonic since $\mathbb{N}\subseteq\mathbb{N}\cup\{\frac{1}{2}\}$, but $W(\mathbb{N})=\mathscr{L}\setminus\mathbb{N}\not\subseteq\mathscr{L}\setminus(\mathbb{N}\cup\{\frac{1}{2}\})=W(\mathbb{N}\cup\{\frac{1}{2}\})$.
    Every logical structure that has an adequate type-I normal $S$-semantics is monotonic by Theorem 4.9. We now show that $(\mathscr{L},W)$, viz., the ones where $W$ satisfies cautious monotonicity or weak cumulative transitivity have adequate type-II normal $S$-semantics.
    Definition 4.19 (Cautious Monotonicity, Weak Cumulative Transitivity).
    Suppose $(\mathscr{L},W)$ is a logical structure.
    1. (i)
      $W$ is said to satisfy cautious monotonicity if for all $\Gamma\cup\Sigma\subseteq\mathscr{L}$, $\Gamma\subseteq\Sigma\subseteq W(\Gamma)$ implies that $W(\Gamma)\subseteq W(\Sigma)$. In this case, $(\mathscr{L},W)$ is said to be of cm-type.
    2. (ii)
      $W$ is said to satisfy weak cumulative transitivity if for all $\Gamma\cup\Sigma\subseteq\mathscr{L}$, $\Gamma\subseteq\Sigma\subseteq W(\Gamma)$ implies that $W(\Sigma)\subseteq W(\Gamma)$. In this case, $(\mathscr{L},W)$ is said to be of wct-type.
    Remark 4.20.
    ‘Cautious monotonicity’ was introduced by Gabbay in 1985 under the name ‘restricted monotoncity’ (see, e.g., [20, p. 455]). In the above definition, we have used Makinson’s version of ‘cautious monotonicity’ instead (see, e.g., [23, p. 43]). The name ‘weak cumulative transitivity’ is taken from [33, p. 230].
    Theorem 4.21 (Suszko Reduction for $cm$-type).
    Let $(\mathscr{L},W)$ be a $cm$-type logical structure and $\mathsf{S}=(\mathbf{B},R_{cm},\models,\mathcal{P}(\mathscr{L}))$ be a normal $S$-semantics for $\mathscr{L}$ satisfying the following properties.
    1. $\bullet$
      $\mathbf{B}=\{\chi_{\Gamma}:\Gamma\subseteq\mathscr{L}\}\cup\{\chi_{W(\Gamma)}:\Gamma\subseteq\mathscr{L}\}$.
    2. $\bullet$
      $R_{cm}=\displaystyle\bigcup_{\Gamma\subseteq\mathscr{L}}R_{\Gamma}$, where for each $\Gamma\subseteq\mathscr{L}$, $R_{\Gamma}$ is such that
      $(\chi_{\Sigma},\chi_{W(\Sigma)})\in R_{\Gamma}~\text{iff}~\Sigma=\Gamma~\text{or}~\Sigma\subseteq W(\Gamma)$
    Then, $W=W^{\mathsf{S}}_{II}$.
    Proof.
    Let $\Gamma\subseteq\mathscr{L}$ such that $W(\Gamma)\not\subseteq W^{\mathsf{S}}_{II}(\Gamma)$. Then, there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W^{\mathsf{S}}_{II}(\Gamma)$. Then, for all $U\subseteq R_{cm}$, $(\chi_{\Gamma},\chi_{W(\Gamma)})\notin U$ or there exists $(\chi_{\Sigma},\chi_{W(\Sigma)})\in U$ such that $\chi_{\Sigma}\models\Gamma$ but $\chi_{W(\Sigma)}\not\models\{\alpha\}$. Now, $(\chi_{\Gamma},\chi_{W(\Gamma)})\in R_{\Gamma}$. So, there exists $\Sigma\subseteq\mathscr{L}$ with $(\chi_{\Sigma},\chi_{W(\Sigma)})\in R_{\Gamma}$ such that $\chi_{\Sigma}\models\Gamma$ but $\chi_{W(\Sigma)}\not\models\{\alpha\}$.
    Now, $\chi_{\Sigma}\models\Gamma$ implies that $\Gamma\subseteq\Sigma$, and $\chi_{W(\Sigma)}\not\models\{\alpha\}$ implies that $\alpha\notin W(\Sigma)$. Since, $\alpha\in W(\Gamma)$, $\Sigma\neq\Gamma$. Then, as $(\chi_{\Sigma},\chi_{W(\Sigma)})\in R_{\Gamma}$, $\Sigma\subseteq W(\Gamma)$. Thus, $\Gamma\subseteq\Sigma\subseteq W(\Gamma)$, and hence by cautious monotonicity, $W(\Gamma)\subseteq W(\Sigma)$. Then, since $\alpha\notin W(\Sigma)$, $\alpha\notin W(\Gamma)$ - a contradiction. Thus, $W(\Gamma)\subseteq W^{\mathsf{S}}_{II}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$.
    Next, let $\Gamma\subseteq\mathscr{L}$ be such that $W^{\mathsf{S}}_{II}(\Gamma)\not\subseteq W(\Gamma)$. So, there exists $\alpha\in W^{\mathsf{S}}_{II}(\Gamma)$ such that $\alpha\notin W(\Gamma)$. Now $R_{\Gamma}\subseteq R_{cm}$ and $(\chi_{\Gamma},\chi_{W(\Gamma)})\in R_{\Gamma}$. Then, since $\chi_{\Gamma}\models\Gamma$ and $\chi_{W(\Gamma)}\not\models\{\alpha\}$, $\alpha\notin W^{\mathsf{S}}_{II}(\Gamma)$ - a contradiction. Thus, $W^{\mathsf{S}}_{II}(\Gamma)\subseteq W(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$. Hence $W=W^{\mathsf{S}}_{II}$. ∎
    Theorem 4.22 (Suszko Reduction for Weak Cumulative Transitive Logical Structures).
    Let $(\mathscr{L},W)$ be a wctwct-type logical structure and $\mathsf{S}=(\mathbf{B},R_{wct},\models,\mathcal{P}(\mathscr{L}))$ be a normal $S$-semantics for $\mathscr{L}$ satisfying the following properties.
    1. $\bullet$
      $\mathbf{B}=\{\chi_{\Gamma}:\Gamma\subseteq\mathscr{L}\}\cup\{\chi_{W(\Gamma)}:\Gamma\subseteq\mathscr{L}\}$.
    2. $\bullet$
      $R_{wct}=\displaystyle\bigcup_{\Gamma\subseteq\mathscr{L}}R_{\Gamma}$ where for each $\Gamma\subseteq\mathscr{L}$,
      $R_{\Gamma}=\{(\chi_{\Gamma},\chi_{W(\Gamma)})\}\cup\{(\chi_{W(\Sigma)},\chi_{W(\Sigma)}):\Sigma\subseteq\Gamma\}$
    Then, $W=W^{\mathsf{S}}_{II}$.
    Proof.
    Let $\Gamma\subseteq\mathscr{L}$ such that $W(\Gamma)\not\subseteq W^{\mathsf{S}}_{II}(\Gamma)$. Then, there exists $\alpha\in W(\Gamma)$ such that $\alpha\notin W^{\mathsf{S}}_{II}(\Gamma)$. Then, for all $U\subseteq R_{wct}$, $(\chi_{\Gamma},\chi_{W(\Gamma)})\notin U$ or there exists $(u,v)\in U$ such that $u\models\Gamma$ but $v\not\models\{\alpha\}$. Since $(\chi_{\Gamma},\chi_{W(\Gamma)})\in R_{\Gamma}$, there exists $(u^{\prime},v^{\prime})\in R_{\Gamma}$ such that $u^{\prime}\models\Gamma$ but $v^{\prime}\not\models\{\alpha\}$. Moreover, $v^{\prime}\not\models\{\alpha\}$ implies that $v^{\prime}\neq\chi_{W(\Gamma)}$, and so $(u^{\prime},v^{\prime})\neq(\chi_{\Gamma},\chi_{W(\Gamma)})$. Therefore, $u^{\prime}=v^{\prime}=W(\Delta)$ for some $\Delta\subseteq\mathscr{L}$. Now, $\chi_{W(\Delta)}\models\Gamma$ implies that $\Gamma\subseteq W(\Delta)$, and $\chi_{W(\Delta)}\not\models\{\alpha\}$ implies that $\alpha\notin W(\Delta)$. Furthermore, since $(\chi_{W(\Delta)},\chi_{W(\Delta)})\in R_{\Gamma}$, $\Delta\subseteq\Gamma$ which implies that $\Delta\subseteq\Gamma\subseteq W(\Delta)$. Therefore, by weak cumulative transitivity, $W(\Gamma)\subseteq W(\Delta)$. Then, as $\alpha\notin W(\Delta)$, $\alpha\notin W(\Gamma)$ - a contradiction. Thus, $W(\Gamma)\subseteq W^{\mathsf{S}}_{II}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$.
    Next, let $\Gamma\subseteq\mathscr{L}$ such that $W^{\mathsf{S}}_{II}(\Gamma)\not\subseteq W(\Gamma)$. So, there exists $\alpha\in W^{\mathsf{S}}_{II}(\Gamma)$ such that $\alpha\notin W(\Gamma)$. Now $R_{\Gamma}\subseteq R_{cm}$ and $(\chi_{\Gamma},\chi_{W(\Gamma)})\in R_{\Gamma}$. Then, since $\chi_{\Gamma}\models\Gamma$ and $\chi_{W(\Gamma)}\not\models\{\alpha\}$, $\alpha\notin W^{\mathsf{S}}_{II}(\Gamma)$ - a contradiction. Thus, $W^{\mathsf{S}}_{II}(\Gamma)\subseteq W(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$. Hence $W=W^{\mathsf{S}}_{II}$. ∎
    5 What is a Many-valued Logical Structure?
    A many-valued logical structure is usually thought to be a logical structure which does not have an adequate bivalent semantics. The term ‘bivalent semantics’ is generally understood via the notion of logical matrix (see, e.g., [26, p. 30]). However, as we have pointed out in §2, there is no formal definition of bivalent semantics. The definition based on logical matrices is problematic, as pointed out in [14, p. 283].
[…]
    This problem is perhaps of even greater significance while attempting to define many-valued logical structures. In fact, one might argue that there is no a priori reason to be ‘confined’ by any particular definition of many-valued logical structures based on a specific notion of semantics, including the ones proposed in Definition 2.1 or Definition 4.1.
    However, if we are willing to maintain such a relativistic perspective, we have to accept that different definitions of semantics will lead to different conclusions regarding $\mathbf{ST}$ (cf. e.g., Theorem 3.10 and Theorem 4.12), and hence, potentially different, mutually incomparable definitions of many-valued logical structures.
    Although this might be mathematically enriching, it is perhaps unsatisfactory from a philosophical perspective because such an approach fails to elucidate the intuition behind ‘many-valuedness’. Consequently, one cannot help but wonder whether something more fundamental is hidden underneath this potential menagerie of notions. The issue of many-valuedness is perhaps intimately connected with the principle of bivalence. We expound on this connection in the rest of this section.
    The following statements are two very general formulations of the principle of bivalence in the context of logical structures.
    $\mathbf{PoB}^{P}_{1}:$ There is a property $P$ such that for every logical structure $(\mathscr{L},W)$, exactly one of the following statements holds:
    1. (a)
      $(\mathscr{L},W)$ satisfies $P$.
    2. (b)
      $(\mathscr{L},W)$ does not satisfy $P$.
    $\mathbf{PoB}^{P}_{2}:$ For every logical structure $(\mathscr{L},W)$, there is a property $P$ such that exactly one of the following statements holds:
    1. (a)
      $(\mathscr{L},W)$ satisfies $P$.
    2. (b)
      $(\mathscr{L},W)$ does not satisfy $P$.
    One cannot help but notice the similarity between these formulations of the principle of bivalence and $\mathbf{ST}$. Consider $\mathbf{PoB}^{P}_{1}$ for example. If one subscribes to the truth-theoretic conception of the notion of logical validity - the view that in logical inferences, ‘truth’ is necessarily preserved from premise(s) to conclusion(s) - and maintains that there is a generic notion of ‘truth’, then $\mathbf{ST}$ can be seen as a principle of bivalence of the form $\mathbf{PoB}^{P}_{1}$. To see this, suppose $P$ is a property and a logical structure $(\mathscr{L},W)$ satisfies $P$ iff it has an adequate bivalent semantics. This means that every element of $\mathscr{L}$ satisfies (semantic) property (not necessarily the same as $P$) or not. In the context of ST, it is ‘designation’, which Suszko identifies with ‘truth’. We quote in this context the following passage from [36, pp. 8-9] (emphasis ours) .
    [T]o identify a property with truth, at the bare minimum, it must be possible to treat such a property as a semantic value of sentences. This is a reasonable assumption, since truth and falsehood are not just semantic properties, but the semantic properties par excellence of sentences. If a property could not be treated as a value of sentences in a suitable semantics, then an identification with truth would be prevented. Suszko’s theorem fills this possible gap between designation and truth by showing that designation can actually be turned into a semantic (namely algebraic) value. Clearly, the mere possibility of treating a property as a semantic referent of a suitable semantic is not sufficient to conclude that it represents truth. After all, a semantic value could also stand for other semantic properties, like, for instance, falsehood. Suszko’s theorem alone only shows that designation can be identified with a semantic value, but it does not show that it is a semantic value for truth. This, however, is a little step easily covered once it is accepted that validity consists in necessary truth preservation. If valid inferences are those necessarily preserving truth, and designation is the value preserved in valid inferences, then designation and truth must be identical. That validity consists in truth preservation is a reasonable philosophical assumption whose legitimacy Suszko accepts without further inquiry, and that it is also assumed in the truth pluralist debate. The crucial and critical step, thus, is exactly the one handled by Suszko’s theorem, which shows that designation can be turned into a semantic (algebraic) value of an adequate semantics.
    This identification, however, is not maintained by every truth pluralist [^7] A truth pluralist is person who believes ‘that sentences belonging to different areas of discourse can be, and in some cases are, true in different ways’ [36, p. 1] even if they believe in the truth-theoretic account of logical validity. The reason behind this is explained succinctly in the following passage from [36, p. 3] (emphasis ours).
[…]
    However, not committing to a truth-theoretic account of validity is not going to help if one is a pluralist relative to some other property $P$ as well. Moreover, not being a pluralist relative to $P$ will require further justification.
    Nevertheless, one can maintain $\mathbf{PoB}^{P}_{2}$ without any qualms since in this case the property $P$ depends on the logical structure under consideration. More precisely, given two properties, $P$ and $Q$, it is possible that a logical structure satisfies $\mathbf{PoB}^{Q}_{2}$ but not for $\mathbf{PoB}^{P}_{2}$. Then, it may be referred to as ‘many-valued relative to $P$’ but ‘two-valued relative to $Q$’.
    We note the following passage from [36, pp. 9-10] (emphasis ours).
    To make the point more vivid, and clearly see why the theorem is crucial to defend the philosophical thesis, suppose that Suszko’s theorem did not go through, and designation could not be turned into an algebraic value of a suitable semantics. In this case, two ingredients would be disentangled in the formulation of a many-valued logic: a level of semantic properties (like truth, falsehood, and possibly others) represented by algebraic values and a level of logical notions employed to define validity (like designation and non-designation). Algebraic values would correspond to properties of sentences that characterise their semantic status. Logical notions, by contrast, would not characterise any semantic feature of sentences, but just display their logical relations. While certainly connected, logic and semantics would be taken apart. In particular, logical notions like designation would not be, even implicitly, semantic values. Designation, far from being a value standing for a semantic property on its own, would just be a meta-logical notion employed to characterise what relations among semantic values take place in valid inferences. In such a case, the philosophical claim that every logic is logically two-valued would vanish. First of all, with the reduction neutralised, the semantic values transmitted in valid inferences could only be those in the many-valued semantics, because there would not be other values at all. Strictly speaking, designation could not be a preserved value, because it could not even be a value. Secondly, given that designation could not be a value, let alone the preserved value, it could not be identified with truth. Thirdly, given that designation would not stand for a value, logical bivalence (reflected in the bivalent distinction between designated and non-designated) would not collapse into semantic two-valuedness. Many-valued logic would be authentically semantically many-valued and truth one of the semantic values—and indeed the preserved value—despite logical validity being defined in bivalent terms by means of bivalent meta-theoretical notions. In this way, the threat to many-valued logic would be thwarted.
    Suszko’s Theorem (Suszko Reduction) thus shows us that the metalogical property of designation can be ‘expressed’ in the object language. The fact that there can only be two possibilities, i.e., designation and non-designation, is thus crucial to Suszko Reduction. However, this dichotomy is nothing but a meta-linguistic assumption about designation. This point has been made, e.g., in [15, p. 390].
    One of the ways to understand Suszko’s distinction between “algebraic values” and “logical values” is to interpret the latter as the truth values of the metatheory, while the former would be the truth values of the theory.
    In other words, ST can be seen as a pointer to the bivalence of the metatheory. Now, what happens if the metatheory is not two-valued, or is it even possible for the metatheory to not be two-valued?
    Caleiro et al. in [10, p. 3] say that “there is some metalinguistic bivalence that one will not easily get rid of: either an inference obtains or it does not, but not both”. However, many-valued consequence relations at the meta-level have been discussed and investigated thoroughly in [12]. A quotation from [12, p. 11] would be relevant here (emphasis ours).
    [T]he objective of the theory of graded consequence (GCT) is to introduce many-valuedness to the consequence relation and other related notions such as consistency, tautologihood, etc. This is in the same direction as had been pursued during the origination and development of many-valued logics. While in many-valued logics the object-level formulae are considered to be many-valued, meta-level notions like consequence are taken as two-valued; in GCT, items of both the levels are treated as many-valued. Various reasons may be ascribed for lifting the meta-level notions to many-valued ones.
    We think that it is possible to give a precise definition of many-valued logical structures with the help of GCT. The rest of this section is devoted to realising one such possibility.
    We begin by observing that the crispness of $\models$ in $S_{\mathfrak{S}}$ relies on the crispness of $\models_{i}$’s (see Theorem 4.4 and Theorem 4.5). We, therefore, propose a notion of many-valued logical structures via many valued $\models_{i}$’s, similar to what has been done in GCT. More specifically, we define the following.
    Definition 5.1 ($\kappa$-valued Semantics of Order $\lambda$).
    Given a set $\mathscr{L}$, a cardinal $\kappa$ and an ordinal $\lambda>0$, a $\kappa$-valued semantics of order $\lambda$ for a family of sets $\{\mathscr{L}_{i}\}_{i\in\lambda}$ is a tuple of the following form $\mathfrak{S}=(\{\mathbf{M}_{i}\}_{i\in\lambda},\{\models_{i}\}_{i\in\lambda},\{A_{i}\}_{i\in\lambda},S,\{\mathcal{P}(\mathscr{L}_{i})\}_{i\in\lambda})$ satisfying the following conditions.
    1. (i)
      $\mathscr{L}_{0}=\mathscr{L}$.
    2. (ii)
      For each $i\in\lambda$, $\mathbf{M}_{i},A_{i}$ are sets, $|A_{i}|\geq\kappa$, and $\models_{i}:\mathbf{M}_{i}\times\mathcal{P}(\mathscr{L}_{i})\to A_{i}$ for each $i\in\lambda$.
    3. (iii)
      There exists $i\in\lambda$ such that $|A_{i}|=\kappa$.
    4. (iv)
      $\emptyset\subsetneq S\subseteq\{((\models_{i},a_{i}),(\models_{j},a_{j})):(a_{i},a_{j})\in A_{i}\times A_{j}\}$.
    Given a $\kappa$-valued semantics of order $\lambda$ for a family of sets $\{\mathscr{L}_{i}\}_{i\in\lambda}$, the entailment relation induced by $\mathfrak{S}$, denoted by $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\kappa,\lambda)}_{\mathfrak{S}}$ is defined as follows. For all $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$,
    $\Gamma\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\kappa,\lambda)}_{\mathfrak{S}}\alpha\iff$ $~\text{for all}~((\models_{i},a_{i}),(\models_{j},a_{j}))\in S~\text{and}~(m_{i},m_{j})\in\mathbf{M}_{i}\times\mathbf{M}_{j},$
    $\models_{i}(m_{i},\Gamma)=a_{i}~\text{implies that}~\models_{j}(m_{j},\{\alpha\})=a_{j}$
    We now define a many-valued logical structure as follows.
    Definition 5.2 ($\kappa$-valued Logical Structure of Order $\lambda$).
    Suppose $\kappa$ is a cardinal, $\lambda>0$ is an ordinal, $\{(\mathscr{L}_{i},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{i})\}_{i\in\lambda}$ is a family of logical structures and $\mathfrak{S}=(\{\mathbf{M}_{i}\}_{i\in\lambda},\{\models_{i}\}_{i\in\lambda},\{A_{i}\}_{i\in\lambda},S,\{\mathcal{P}(\mathscr{L}_{i})\}_{i\in\lambda})$ is a $\kappa$-valued semantics of order $\lambda$. A logical structure $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$ is said to be $\kappa$-valued of order $\lambda$ for $\{(\mathscr{L}_{i},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{i})\}_{i\in\lambda}$ relative to $\mathfrak{S}$ if the following conditions hold.
    1. (i)
      For all $i,j\in\lambda$, with $i\in j$, there exists an injective function from $\mathscr{L}_{j}$ to $\mathscr{L}_{i}$.
    2. (ii)
      $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,=\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\kappa,\lambda)}_{\mathfrak{S}}$.
    3. (iii)
      For every $\mu$-valued semantics $\mathfrak{T}=(\{\mathbf{N}_{i}\}_{i\in\lambda},\{\models_{i}\}_{i\in\lambda},\{A_{i}\}_{i\in\lambda},S,\{\mathcal{P}(\mathscr{L}_{i})\}_{i\in\lambda})$ of order $\lambda$ for $\{\mathscr{L}_{i}\}_{i\in\lambda}$, with $\mu<\kappa$, $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,\neq\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\mu,\lambda)}_{\mathfrak{S}}$.
    A logical structure $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$ is said to be $\kappa$-valued of order $\lambda$, if there exists a semantics $\mathfrak{S}$ and a family of sets $\{\mathscr{L}_{i}\}_{i\in\lambda}$ such that $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$ is $\kappa$-valued of order $\lambda$ for $\{\mathscr{L}_{i}\}_{i\in\lambda}$ relative to $\mathfrak{S}$.
    Remark 5.3.
    It is imperative to emphasise that even though we defined of $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\kappa,\lambda)}_{\mathfrak{S}}$ in a specific manner, other definitions are possible. Thus, a definition of $\kappa$-valued logical structure depends on how $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\kappa,\lambda)}_{\mathfrak{S}}$ is defined.
    Remark 5.4.
    1. (a)
      The key idea behind the above definition is the observation of the distinction between object and metalanguage and the resulting hierarchy of languages (see, e.g., [22]). This the main reason for considering a family of logical structures $\{(\mathscr{L}_{i},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{i})\}_{i\in\lambda}$. This family is indexed via ordinals so as to introduce an ordering among the $\mathscr{L}_{i}$’s in such a way that for all $i\in\lambda$, $\mathscr{L}_{i+1}$ can be thought of as a metalanguage of $\mathscr{L}_{i}$.
    2. (b)
      To distinguish between the use and mention of a word or symbol, we sometimes put quotation marks around that symbol whenever we mention it. In the context of (formal) object language and (formal) metalanguage, this means that for every formula $\alpha$ of the object language, $\alpha$’ is a constant symbol of the metalanguage (see, [13, Footnote 136] and [12, Section 6.2]). This gives rise to a natural injective function from the metalanguage to the object language under consideration. Property (i) of Definition 5.2 has been formulated with this issue in mind.
    3. (c)
      Each $\mathbf{M}_{i}$ can be thought of as a ‘class of models’ of the language $\mathscr{L}_{i}$. Similarly, $\models_{i}$’s can be thought of as providing the ‘satisfaction relation’ between the models and the corresponding languages.
    4. (d)
      The motivation for (iii) and (iv) are similar to that of Remark 2.10.
    Since this hierarchy of languages is not unique for a given (object) language, the crux of all these discussions is that many-valuedness of a logical structure is not dependent on the logical structure alone but also on the language/meta-language hierarchy. Hence, interpreting ‘logical value’ as truth-values of the metatheory/metalogic makes us wonder about the reason for ignoring the truth-values of the higher order metatheories/metalogics. A justification for this, however, can be provided by observing that in most cases, it suffices to consider the meta-metalogic to be a fragment of classical propositional logic, as e.g., in [12]. Nevertheless, there is no a priori reason for not considering a many-valued meta-metalogic, and continuing in this line of argument, we can conclude that there is no reason to consider any one of the higher metalogic to be two-valued.
    This motivates us to propose the following generalisations of $\mathbf{ST}$.
    1. (a)
      Let $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$ be a logical structure which is $\kappa$-valued of order $\lambda$ for $\{(\mathscr{L}_{i},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{i})\}_{i\in\lambda}$ relative to a $\kappa$-valued semantics of order $\lambda$. Then, there exists $i\in\lambda$ and an ordinal $\gamma$ such that $(\mathscr{L}_{i},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{i})$ is $2$-valued of order $\gamma$.
    2. (b)
      Let $(\mathscr{L},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-})$ be a logical structure which is $\kappa$-valued of order $\lambda$ for $\{(\mathscr{L}_{i},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{i})\}_{i\in\lambda}$ relative to a semantics $\kappa$-valued semantics of order $\lambda$. Then, for all $i\in\lambda$ there exists a cardinal $\sigma_{i}$ and an ordinal $\gamma_{i}$ with $\sigma_{i}\leq\kappa$ such that $(\mathscr{L}_{i},\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}_{i})$ is $\sigma_{i}$-valued of order $\gamma_{i}$.
    6 Concluding Remarks
    The main contributions of this article to the study of $\mathbf{ST}$, in particular, and to the study of many-valued logics, in general, are as follows.
    1. $\bullet$
      A rather generalised notion of semantics has been proposed. We have also showed that every logical structure has a canonical adequate semantics (Theorem 2.6).
    2. $\bullet$
      A precise definition of inferentially $\kappa$-valued logical structures has been given. We have also proved that (a) the logical structures induced by $q$-, $p$-, $s_{\kappa}$-consequence operators are at least inferentially $3$-valued (see Theorem 3.10, Theorem 3.17 and Theorem 3.33 for details), and (b) the $r_{k}$-consequence operators are inferentially $4$-valued (Theorem 3.40).
    3. $\bullet$
      We have generalised Suszko Reduction, introduced the notion of Suszko semantics, and based upon this notion, showed that several logical structures have an adequate bivalent semantics (see §4). The constructions of adequate bivalent semantics for cautious monotonic and weak cumulative transitive logical structures are, according to us, the most important contributions of this section.
    4. $\bullet$
      An explicit definition of $\kappa$-valued logical structure’ (Definition 5.2) and generalised versions of $\mathbf{ST}$ are discussed.
    There are, however, several questions which need to be investigated further. These are left for future work. We list some of them below.
    1. $\bullet$
      Is it possible to obtain an adequate $S$-semantics for every non-monotonic logical structure, perhaps by changing the notion of $S$-entailment?
    2. $\bullet$
      Given an ordinal $\lambda$ and $n\in\mathbb{N}$, what exactly are the distinguishing properties between a $n$- and a $n+1$-valued logical structure or order $\lambda$?
    3. $\bullet$
      Is it possible to find a minimal adequate $S$-semantics for every logical structure which admits an adequate $S$-semantics?
    4. $\bullet$
      Is it possible to give example(s) of logic(s) for which both the generalisation of Suszko’s Thesis, as proposed above, fails?